### v1

# On automorphism groups of smooth hypersurfaces

Abstract.
We show that smooth hypersurfaces in complex projective spaces with automorphism groups of maximum size are isomorphic to Fermat hypersurfaces, with a few exceptions. For the exceptions, we give explicitly the defining equations and automorphism groups.
1. Introduction
The purpose of this paper is to study smooth hypersurfaces in complex projective spaces with large automorphism groups. Let $F$ in ${\mathbb{C}}[x_{1},x_{2},\ldots,x_{n+2}]$ be a smooth homogeneous polynomial of degree $d\geq 3$ and let $X_{F}\subset{\mathbb{P}}^{n+1}$ denote the smooth hypersurface defined by $F$, where $n\geq 1$. An automorphism $f\in{\rm Aut}(X_{F})$ of $X_{F}$ is called linear with respect to the embedding $X_{F}\hookrightarrow{\mathbb{P}}^{n+1}$ if $f$ extends to an automorphism of ${\mathbb{P}}^{n+1}$, i.e., $f$ is given by a linear change of homogeneous coordinates. The automorphism group ${\rm Aut}(X_{F})$ is finite and equal to its subgroup ${\rm Lin}(X_{F})$ of linear automorphisms if $(n,d)\neq(1,3),(2,4)$ (see [MM63] for $n\geq 2$ and [Cha78] for $n=1$). See Section 2 for some notations on groups. We use ${\rm Aut}(F)$ to denote the group of linear transformations that preserve $F$. Then clearly $\mu_{n+2,d}$ is a normal subgroup of ${\rm Aut}(F)$ and the following exact sequence gives a basic relation between ${\rm Lin}(X_{F})$ and ${\rm Aut}(F)$:
$1\rightarrow\mu_{n+2,d}\longrightarrow{\rm Aut}(F)\longrightarrow{\rm Lin}(X_{F})\rightarrow 1$ (1.1)
Bounding the size of ${\rm Aut}(X_{F})$, ${\rm Lin}(X_{F})$ or ${\rm Aut}(F)$ is almost the same thing and has a long history (see [OS78, Section 6] for some historical remarks). Bott–Tate and Orlik–Solomon showed that there exists an upper bound for the order $|{\rm Aut}(F)|$ depending only on $n$ and $d$ (see [OS78, Corollary (2.7)]). Finding an effective bound in terms of $n$ and $d$ is important. We use $X_{d}^{n}$ to denote the Fermat hypersurface in ${\mathbb{P}}^{n+1}$ of degree $d$ defined by Fermat polynomial $F_{d}^{n}:=x_{1}^{d}+x_{2}^{d}+\cdots+x_{n+2}^{d}$. It is known that ${\rm Aut}(X_{d}^{n})$ is isomorphic to $C_{d}^{n+1}\rtimes S_{n+2}$ and $|{\rm Aut}(X_{d}^{n})|=d^{n+1}\cdot(n+2)!$ if $(n,d)\neq(1,3),(2,4)$ (see [HS80], [Shi88]). By a classical theorem of Hurwitz, the automorphism group of a complex curve of genus $g\geq 2$ is of order at most $84(g-1)$. This implies that for a plane curve $X_{F}$ with $d\geq 4$, $|{\rm Aut}(X_{F})|\leq 42(d-3)d$. Automorphism groups of plane curves have been extensively studied by many people. For $n=1$ and $d\geq 4$, plane curves $X_{F}$ with maximum $|{\rm Aut}(X_{F})|$ are isomorphic to Fermat curves $X_{d}^{1}$ (resp. Klein quartic, resp. Wiman sextic) if $d\neq 4,6$ (resp. $d=4$, resp. $d=6$) (see [Pam13], [Har19] and references therein). For $n\geq 2$, Howard–Sommese [HS80, Theorem 2] proved that there is a constant $k_{n}$ depending only on $n$ such that $|{\rm Aut}(X_{F})|\leq k_{n}\cdot d^{n+1}$ if $(n,d)\neq(2,4)$. Xiao ([Xia94], [Xia95]) proved that a minimal smooth projective surface $S$ of general type has at most $42^{2}K_{S}^{2}$ automorphisms. Consequently, for a surface $X_{F}\subset{\mathbb{P}}^{3}$ with $d\geq 5$, $|{\rm Aut}(X_{F})|\leq 42^{2}(d-4)^{2}d$. For $2\leq n\leq 5$, Fermat cubic $n$-fold has the largest possible order for the automorphism group among all smooth cubic $n$-folds (see [Seg42], [Hos97], [Dol12] for $n=2$; [WY20] for $n=3$; [LZ22], [YYZ23] for $n=4$; [YYZ23] for $n=5$). Fermat quintic $3$-fold has the largest possible order for the automorphism group among all smooth quintic $3$-folds ([OY19]). For some partial results on abelian subgroups of automorphism groups of smooth hypersurfaces of arbitrary dimension, see [CS95], [GL11], [GL13], [Zhe22], [GLM23], etc. However, to the best of our knowledge, for $n\geq 2$ except $(n,d)=(2,3),(2,4),(3,3),(4,3),(5,3)$ and $(3,5)$, the maximum order $|{\rm Aut}(X_{F})|$ is still unknown (see e.g. [Huy23, Remark 3.20], [GLMV24, Remark 2.11], [Ess24, Question 3.9]) and the classification of groups of linear automorphisms of hypersurfaces is in a rudimentary state as also mentioned in [Dol19, Page 6]. Our main result is to completely classify smooth hypersurfaces with automorphism groups of maximum order, which in particular gives an optimal upper bound for $|{\rm Aut}(X_{F})|$.
Theorem 1.1.
Fix integers $n\geq 1$, $d\geq 3$ with $(n,d)\neq(1,3),(2,4)$. Let $X\subset{\mathbb{P}}^{n+1}$ be a smooth hypersurface of degree $d$ with maximum $|{\rm Aut}(X)|$. Then
$|{\rm Aut}(X)|=d^{n+1}\cdot(n+2)!\;\text{ and }\;X\text{ is isomorphic to Fermat hypersurface }X_{d}^{n}$
with the following exceptions:
$(n,d)$ ${\rm Aut}(X)$ $|{\rm Aut}(X)|$ $X\text{ is isomorphic to }X_{F}$
$(1,4)$ ${\rm PSL}(2,7)$ $168$ $F=x_{1}^{3}x_{2}+x_{2}^{3}x_{3}+x_{3}^{3}x_{1}$
$(1,6)$ $A_{6}$ $360$ $F=10x_{1}^{3}x_{2}^{3}+9(x_{1}^{5}+x_{2}^{5})x_{3}-45x_{1}^{2}x_{2}^{2}x_{3}^{2}-135x_{1}x_{2}x_{3}^{4}+27x_{3}^{6}$
$(2,6)$ $C_{6}.(S_{4}^{2}\rtimes C_{2})$ $6912$ $F=x_{1}^{5}x_{2}+x_{2}^{5}x_{1}+x_{3}^{5}x_{4}+x_{4}^{5}x_{3}$
$(2,12)$ $C_{12}.(A_{5}^{2}\rtimes C_{2})$ $86400$ $F=x_{1}^{11}x_{2}+11x_{1}^{6}x_{2}^{6}-x_{1}x_{2}^{11}+x_{3}^{11}x_{4}+11x_{3}^{6}x_{4}^{6}-x_{3}x_{4}^{11}$
$(4,6)$ ${\rm PSU}(4,3).C_{2}$ $6531840$ $F=\sum\limits_{1\leq i\leq 6}x_{i}^{6}+\sum\limits_{1\leq i\neq j\leq 6}15x_{i}^{4}x_{j}^{2}-\sum\limits_{1\leq i<j<k\leq 6}30x_{i}^{2}x_{j}^{2}x_{k}^{2}+240\sqrt{-3}x_{1}x_{2}x_{3}x_{4}x_{5}x_{6}$
$(4,12)$ $C_{12}^{2}.(A_{5}^{3}\rtimes S_{3})$ $186624000$ $F=x_{1}^{11}x_{2}+11x_{1}^{6}x_{2}^{6}-x_{1}x_{2}^{11}+x_{3}^{11}x_{4}+11x_{3}^{6}x_{4}^{6}-x_{3}x_{4}^{11}+x_{5}^{11}x_{6}+11x_{5}^{6}x_{6}^{6}-x_{5}x_{6}^{11}$
The sextic $F$ and its automorphism group in the case $(n,d)=(4,6)$ were given by Todd [Tod50]. It seems that $X_{F}$ in the cases $(n,d)=(2,6),(2,12),(4,12)$ in the table above are previously unknown. It is known that automorphism groups of smooth hypersurfaces of general type are equal to their birational automorphism groups. Hacon–${\rm M^{c}Kernan}$–Xu [HMX13] showed that the number of birational automorphisms of a variety of general type $X$ is bounded by $c\cdot{\rm vol}(X,K_{X})$, where $c$ is a constant that only depends on the dimension of $X$, and they asked for finding an explicit bound for the constant $c$ ([HMX13, Question 1.2]). As a direct consequence of Theorem 1.1, we obtain the optimal values of such constants for smooth hypersurfaces of general type.
Corollary 1.2.
Let $X\subset{\mathbb{P}}^{n+1}$ be a smooth hypersurface of general type, where $n\geq 2$. Then
$|{\rm Aut}(X)|\leq(n+2)!(n+3)^{n}\cdot{\rm vol}(X,K_{X}).$
Moreover, the equality holds if and only if $X$ is isomorphic to Fermat hypersurface $X_{n+3}^{n}$.
Next we briefly explain ideas of the proof of Theorem 1.1. In fact, we completely classify smooth hypersurfaces $X_{F}$ with $|{\rm Aut}(X_{F})|\geq|{\rm Aut}(X_{d}^{n})|$ (Theorem 7.5), which clearly implies Theorem 1.1. In order to prove Theorem 7.5, by the exact sequence (1.1), it suffices to classify $X_{F}$ with $|{\rm Aut}(F)|\geq|{\rm Aut}(F_{d}^{n})|$. A famous theorem of Jordan says that, for every integer $r\geq 1$, there exists a constant depending only on $r$ such that any finite subgroup of the general linear group ${\rm GL}(r,{\mathbb{C}})$ has an abelian normal subgroup of index at most this constant. Collins [Col07] gave the optimal values ${\rm J}(r)$ of such constants for all $r\geq 1$. By Jordan’s Theorem, as a finite linear group in ${\rm GL}(n+2,{\mathbb{C}})$, ${\rm Aut}(F)$ has a normal abelian subgroup $N$ of index at most ${\rm J}(n+2)$, and $|{\rm Aut}(F)|$ is bounded by the product $|N|\cdot{\rm J}(n+2)$. Hence bounding the size of $N$ leads to an upper bound of $|{\rm Aut}(F)|$. The known bounds for $|{\rm Aut}(F)|$ or $|{\rm Aut}(X_{F})|$ mentioned previously are obtained in this way or a similar way (see [OS78], [HS80]). However, it is often that when the index $[{\rm Aut}(F):N]$ gets bigger, the size of $N$ gets smaller. Thus the product of ${\rm J}(n+2)$ and the largest possible value for $|N|$ rarely gives an optimal bound for $|{\rm Aut}(F)|$ (see Remark 7.6). To overcome this difficulty, we introduce two new notions, canonical bound (Definition 3.5) and Fermat-test ratio (Definition 4.5). Roughly speaking, canonical bounds integrate bounding abelian subgroups of ${\rm Aut}(F)$ with bounding certain primitive groups called primitive constituents (see $\bf{Set\textrm{-}up}$ 3.1) of ${\rm Aut}(F)$, which turns out to be quite effective in our study of bounding $|{\rm Aut}(F)|$. More precisely, based on complete reducibility of linear representations of finite groups and the classical notion of primitive (projective) linear groups (see Section 3), we introduce two associated exact sequences (3.1)-(3.2) for any finite linear group. Combining the associated exact sequences with our bound for abelian subgroups of ${\rm Aut}(F)$ (Theorem 3.4), we show that $|{\rm Aut}(F)|$ is at most the canonical bound $B({\rm Aut}(F))$ of ${\rm Aut}(F)$ (Lemma 3.6). From this, we immediately get an optimal bound for $|{\rm Aut}(F)|$ in the case ${\rm Aut}(F)$ consisting of semi-permutations (Theorem 3.9).
To handle the case ${\rm Aut}(F)$ consisting of not only semi-permutations, we use both canonical bounds and Fermat-test ratios. As a consequence of Jordan’s theorem, for each integer $r\geq 1$, there are only finitely many finite primitive groups in ${\rm PGL}(r,{\mathbb{C}})$ (up to conjugation), and Collins [Col08a, Theorem A] determined the maximum order ${\rm JC}(r)$ among such groups using the well-known classification of finite simple groups. Based on the definition of Fermat-test ratio and the explicit values of ${\rm JC}(r)$, we show that Fermat-test ratios have many nice properties (see Section 4) which are crucial for our classification. In particular, using Fermat-test ratios, we quickly prove that Theorem 7.5 holds if $n\geq 26$ or $d\geq 18$ (Theorem 4.13). Sections 5, 6 and 7 are devoted to the proof of Theorem 7.5 for the remaining (finitely many) pairs $(n,d)$ satisfying $n\leq 25$ and $d\leq 17$. For such pairs of $(n,d)$, we classify $X_{F}$ with $|{\rm Aut}(F)|\geq|{\rm Aut}(F_{d}^{n})|$ in three steps:
  1. (1)
    ${\rm Aut}(F)$ are primitive linear groups (Theorem 6.1);
  2. (2)
    ${\rm Aut}(F)$ are imprimitive linear groups (Theorem 7.1);
  3. (3)
    ${\rm Aut}(F)$ are reducible linear groups (Theorem 7.4).
In these steps, our classification is based on known classifications of finite primitive linear groups in small degrees (see Section 6) and close relations among smoothness and shape of the defining polynomial $F$, canonical bounds, and Fermat-test ratios (see Section 5). Especially, we show that monomials in $F$ involving different primitive constituents of ${\rm Aut}(F)$ considerably reduce the order $|{\rm Aut}(F)|$ (Lemmas 5.4, 5.5, 5.6, 5.7), which plays a key role in steps (2) and (3).
We conclude the introduction by posing a question closely related to our work. Collins [Col08b] obtained the optimal values of modular analogues of Jordan constants for finite linear groups. Dolgachev–Duncan [DD19] classified all possible automorphism groups of smooth cubic surfaces over an algebraically closed field of arbitrary characteristic. It would be interesting to study an analogue of Theorem 1.1 in positive characteristic.
Question 1.3.
Let $k$ be an algebraically closed field of characteristic $p>0$ and let integers $n\geq 1$, $d\geq 3$ with $(n,d)\neq(1,3),(2,3),(2,4)$. Classify smooth hypersurfaces $X\subset{\mathbb{P}}_{k}^{n+1}$ of degree $d$ with maximum $|{\rm Aut}(X)|$.
Acknowledgements
We would like to thank Professors Keiji Oguiso and Zhiwei Zheng for valuable conversations. This work is partially supported by the National Natural Science Foundation of China (No. 12171351, No. 12071337, No. 11921001).
2. Notation and conventions
(2.1)  Let $W$ be a complex vector space of dimension $r\geq 1$. Let $W^{*}$ denote the dual vector space of $W$ and $S(W^{*})$ be the symmetric algebra of $W^{*}$. Each element in $S(W^{*})$ can be viewed as a complex-valued function on $W$. Then ${\rm GL}(W)$ acts from left (resp. right) on $W$ (resp. $S(W^{*})$) via $(f,w)\mapsto f(w)$ (resp. $(F,f)\mapsto f(F)$). Here $f(F)$ is given by $f(F)(w)=F(f(w))$ for any $w\in W$. Note that the action of ${\rm GL}(W)$ on $W$ induces an action of ${\rm GL}(W)$ on ${\mathbb{P}}(W)$ given by $(f,[w])\mapsto[f(w)]$. Two elements $F,F^{\prime}$ in $S(W^{*})$ are called isomorphic to each other if they lie in the same orbit under the action by ${\rm GL}(W)$ on $S(W^{*})$.
(2.2)  If we choose a basis $(e_{1},\ldots,e_{r})$ of $W$, we have the dual basis $(e_{1}^{*},\ldots,e_{r}^{*})$ of $W^{*}$. Then elements $f\in{\rm GL}(W)$ (resp. $w\in W$, resp. $F\in S(W^{*})$) may be naturally viewed as matrices $A=(a_{ij})\in{\rm GL}(r,{\mathbb{C}})$ (resp. vectors $(w_{1},\ldots,w_{r})\in{\mathbb{C}}^{r}$, resp. polynomials $F(x_{1},\ldots,x_{r})\in{\mathbb{C}}[x_{1},\ldots,x_{r}]$). Under such identifications, the actions of ${\rm GL}(r,{\mathbb{C}})$ on ${\mathbb{C}}^{r}$ and ${\mathbb{C}}[x_{1},\ldots,x_{r}]$ mentioned above are given by
$(A,(w_{1},\ldots,w_{r}))\mapsto(\sum_{i=1}^{r}a_{1i}w_{i},\ldots,\sum_{i=1}^{r}a_{ri}w_{i})\,\,\text{ and }\,\,(F,A)\mapsto A(F):=F(\sum_{i=1}^{r}a_{1i}x_{i},\ldots,\sum_{i=1}^{r}a_{ri}x_{i})$
respectively. We call $(e_{1},\ldots,e_{r})$ (resp. $(e_{1}^{*},\ldots,e_{r}^{*})$) the underlying basis of ${\mathbb{C}}^{r}\cong W$ (resp. ${{\mathbb{C}}^{r}}^{*}\cong W^{*}$). In particular, in this paper, we always assume that the underlying bases of $W$ and $W^{*}$ are dual to each other, and if we choose a new basis $(e_{i}^{\prime})$ as the underlying basis of $W$, it is understood that the underlying basis of $W^{*}$ is changed to the dual basis $({e_{i}^{\prime}}^{*})$.
(2.3)  We say the underlying basis $(e_{i})$ of $W$ is compatible with a decomposition $W=W_{1}\oplus\cdots\oplus W_{s}$ if $e_{n_{1}+\cdots+n_{i-1}+j}\in W_{i}$ for $1\leq i\leq s$, $1\leq j\leq n_{i}$, where $n_{i}={\rm dim}(W_{i})$. In this case, for any integer $d\geq 0$, the space of $d$-forms $S^{d}(W^{*})\subset S(W^{*})$ admits a decomposition
$S^{d}(W^{*})=\mathop{\vbox{\hbox{{$\bigoplus$}}}}_{k_{1}+k_{2}+\cdots+k_{s}=d}S^{k_{1}}(W_{1}^{*})\otimes S^{k_{2}}(W_{2}^{*})\otimes\cdots\otimes S^{k_{s}}(W_{s}^{*}),$
which induces a decomposition
$F=\sum_{k_{1}+k_{2}+\cdots+k_{s}=d}F^{(k_{1},k_{2},\ldots,k_{s})}$
for any $d$-form $F\in S^{d}(W^{*})$. We call $F^{(k_{1},k_{2},\ldots,k_{s})}$ the $S^{k_{1}}(W_{1}^{*})\otimes S^{k_{2}}(W_{2}^{*})\otimes\cdots\otimes S^{k_{s}}(W_{s}^{*})$-component of $F$.
(2.4)  Note that if $A_{1},A_{2}\in{\rm GL}(r,{\mathbb{C}})$, we have $(A_{1}A_{2})(F)=A_{2}(A_{1}(F))$. The automorphism group ${\rm Aut}(F)$ of a homogeneous polynomial $F\in{\mathbb{C}}[x_{1},\ldots,x_{r}]$ of degree $d\geq 1$ is given by
${\rm Aut}(F):=\{A\in{\rm GL}(r,{\mathbb{C}})\mid\,A(F)=F\}.$
We say $F$ is smooth if $(0,\ldots,0)$ is the only solution for $\frac{\partial F}{\partial x_{1}}=\cdots=\frac{\partial F}{\partial x_{r}}=0$. In particular, if $r\geq 3$, a smooth form $F$ defines a smooth hypersurface $X_{F}\subset{\mathbb{P}}^{r-1}$ of dimension $r-2$. We denote by $\pi:{\rm GL}(r,{\mathbb{C}})\rightarrow{\rm PGL}(r,{\mathbb{C}})$ the projection map. For a subgroup $H$ of ${\rm GL}(r,{\mathbb{C}})$ (resp. ${\rm PGL}(r,{\mathbb{C}})$), we say $H$ preserves $F$ if $H=G$ (resp. $H=\pi(G)$) for some subgroup $G\subset{\rm Aut}(F)$. In this case, we call $F$ an $H$-invariant form of degree $d$. We say a monomial $\mathfrak{m}$ of degree $d$ is in $F$ (or $\mathfrak{m}\in F$) if the coefficient of $\mathfrak{m}$ is not zero in the expression of $F$. As in [OY19], for a finite subgroup $H_{1}\subset{\rm PGL}(r,{\mathbb{C}})$, we say $H_{1}$ admits a lifting (resp. an $F$-lifting) if there exists $\widetilde{H}_{1}\subset{\rm GL}(r,{\mathbb{C}})$ (resp. $\widetilde{H}_{1}\subset{\rm Aut}(F)$ ) with $\widetilde{H}_{1}\cong\pi(\widetilde{H}_{1})=H_{1}$.
(2.5)  If $F_{1},\ldots,F_{k}\in{\mathbb{C}}[x_{1},\ldots,x_{r}]$ and $a>0$, we use $(F_{1},\ldots,F_{k})^{a}\cdot{\mathbb{C}}[x_{1},\ldots,x_{r}]$ to denote the $a$-th power of the ideal in ${\mathbb{C}}[x_{1},\ldots,x_{r}]$ generated by $F_{1},\ldots,F_{k}$. For a group $G_{1}$, we denote by $Z(G_{1})$ its center. We use $N.H$ to denote a finite group which fits in a short exact sequence $1\rightarrow N\rightarrow N.H\rightarrow H\rightarrow 1$ Some symbols frequently used in this paper are as follows:
$\xi_{k}$ the $k$-th primitive root $e^{\frac{2\pi i}{k}}$ of unity, where $k$ is a positive integer;
$I_{n}$ the identity matrix of rank $n$;
$\mu_{r,d}$ the subgroup of ${\rm GL}(r,{\mathbb{C}})$ of order $d$ generated by the scalar matrix $\xi_{d}I_{r}$;
$C_{n}$ the cyclic group of order $n$;
$S_{n}$ the symmetric group of degree $n$;
$A_{n}$ the alternating group of degree $n$;
$Q_{n}$ the quaternion group of order $n$.
3. Abelian subgroups and canonical bounds
In this section, we introduce two associated exact sequences (3.1)-(3.2) for any finite linear group using primitive groups and we show an upper bound for abelian subgroups of automorphism groups ${\rm Aut}(F)$ of smooth forms $F$ (Theorem 3.4). Then we introduce the notion of canonical bound for $|{\rm Aut}(F)|$ (Definition 3.5 and Lemma 3.6), which enables us to prove that Fermat forms have the largest automorphism groups among smooth forms with automorphism groups consisting of semi-permutations (Theorem 3.9).
3.1. Primitive constituents and associated exact sequences
Let $F\in{\mathbb{C}}[x_{1},\ldots,x_{r}]$ be a non-zero homogeneous polynomial of degree $d\geq 3$, where $r\geq 1$. The automorphism group ${\rm Aut}(F)$ is finite but its group structure is in general complicated. In order to study ${\rm Aut}(F)$, we introduce some notions for finite linear groups. First, we recall the notion of primitive linear groups (for more details, see e.g., [Bli17, Col07, Col08a]). Let $\rho:G\rightarrow{\rm GL}(r,{\mathbb{C}})$ be an irreducible linear representation of a finite group $G$. We say $\rho$ is primitive (resp. imprimitive) if ${\mathbb{C}}^{r}$ cannot (resp. can) be decomposed as a direct sum of proper subspaces permuted under the action of $G$. A finite subgroup in ${\rm GL}(r,{\mathbb{C}})$ is called primitive (resp. irreducible) if the underlying representation is primitive (resp. irreducible). A normal abelian subgroup of a finite primitive linear group $G$ is contained in the center $Z(G)$ of $G$ (cf. [Col08a, Lemma 1]). We call a finite subgroup $H\subset{\rm PGL}(r,{\mathbb{C}})$ primitive if $H=\pi(\widetilde{H})$ for some finite primitive subgroup $\widetilde{H}\subset{\rm GL}(r,{\mathbb{C}})$. Since the center $Z(\widetilde{H})\subset\widetilde{H}$ is a subgroup of $Z({\rm GL}(r,{\mathbb{C}}))=\{\lambda I_{r}\mid\lambda\in{\mathbb{C}}^{\times}\}$, we have $H\cong\widetilde{H}/Z(\widetilde{H})$.
Using primitive groups, we introduce two exact sequences associated to finite linear groups. For this purpose and later use, we fix some notations.
Set-up 3.1.
Let $G$ be a finite subgroup of ${\rm GL}(r,{\mathbb{C}})$ with $r\geq 1$. Then ${\mathbb{C}}^{r}$ admits a decomposition ${\mathbb{C}}^{r}=V_{1}\oplus\cdots\oplus V_{m}$ as a direct sum of irreducible $G$-stable subspaces for some integer $m\geq 1$. For each $1\leq i\leq m$, there is an integer $k_{i}\geq 1$ and a decomposition $V_{i}=W_{i1}\oplus\cdots\oplus W_{ik_{i}}$ as a direct sum of subspaces permuted (transitively) under the action of $G$ such that
${\rm Stab}_{G}(W_{ij}):=\{g\in G\mid\,g(W_{ij})=W_{ij}\}$
acts primitively (but not necessarily faithfully) on $W_{ij}$. Such permutations form transitive subgroups $K_{i}$ ($i=1,\ldots,m$) of $S_{k_{i}}$, and there is a natural group homomorphism $G\rightarrow K_{i}$. Let $s:=k_{1}+\cdots+k_{m}$. The dimensions $r_{ij}$ of the subspaces $W_{ij}$ are called subdegrees of $G$. Let
$H_{ij}:=\{[h]\mid\,h=g|_{W_{ij}},g\in{\rm Stab}_{G}(W_{ij})\}\subset{\rm PGL}(W_{ij}).$
We call $H_{ij}$ the primitive constituents of $G$ (of degree $r_{ij}$) belonging to $V_{i}$.
Definition 3.2.
Under $\bf{Set\textrm{-}up}$ 3.1, we call the intersection
$P:=\bigcap_{1\leq i\leq m,1\leq j\leq k_{i}}{\rm Stab}_{G}(W_{ij})$
the principal subgroup of $G$. Then we call the following exact sequence the first associated exact sequence of $G$
$1\longrightarrow P\longrightarrow G\longrightarrow K_{1}\times\cdots\times K_{m}.$ (3.1)
Let $\widetilde{H}_{ij}:=\{h\mid\,h=g|_{W_{ij}},g\in{\rm Stab}_{G}(W_{ij})\}$. Then we have surjective group homomorphisms $\phi_{ij}:{\rm Stab}_{G}(W_{ij})\longrightarrow\widetilde{H}_{ij}$, $g\mapsto g|W_{ij}$ and $\pi_{ij}:\widetilde{H}_{ij}\longrightarrow H_{ij}$, $h\mapsto[h].$ Thus, $\widetilde{H}_{ij}$ and $H_{ij}$ are primitive subgroups of ${\rm GL}(W_{ij})$ and ${\rm PGL}(W_{ij})$ respectively. Consider the group homomorphism
$\phi:P\longrightarrow H_{11}\times\cdots\times H_{1k_{1}}\times\cdots\times H_{mk_{m}},\;\;g\mapsto([g|W_{11}],\ldots,[g|W_{1k_{1}}],\ldots,[g|W_{mk_{m}}]).$
Let $N:=\ker(\phi)$. Then we call the following exact sequence the second associated exact sequence of $G$
$1\longrightarrow N\longrightarrow P\longrightarrow H_{11}\times\cdots\times H_{1k_{1}}\times\cdots\times H_{mk_{m}}.$ (3.2)
Clearly, a finite linear group determines uniquely its first and second associated exact sequences up to isomorphisms of exact sequences. After choosing basis for all $W_{ij}$, we obtain a basis of ${\mathbb{C}}^{r}$ and we may identity ${\rm GL}(W_{ij})$ (resp. ${\rm PGL}(W_{ij})$) with ${\rm GL}(r_{ij},{\mathbb{C}})$ (resp. ${\rm PGL}(r_{ij},{\mathbb{C}})$). Then each element $A$ in $P$ can be viewed as a block diagonal matrix $A={\rm diag}(A_{11},\ldots,A_{mk_{m}})$ with blocks $A_{ij}\in{\rm GL}(r_{ij},{\mathbb{C}})$. Note that $N$ is abelian and its elements are of the form ${\rm diag}(\lambda_{11}I_{r_{11}},\lambda_{12}I_{r_{12}},\ldots,\lambda_{mk_{m}}I_{r_{mk_{m}}})$, where $\lambda_{ij}$ are non-zero complex numbers.
Definition 3.3.
Under $\bf{Set\textrm{-}up}$ 3.1, let $l(G):=(r_{1}^{\prime},\ldots,r_{s}^{\prime})$ be the $s$-tuple of positive integers such that $r_{1}^{\prime}\geq r_{2}^{\prime}\geq\cdots\geq r_{s}^{\prime}$ and there exists a permutation $\sigma\in S_{s}$ satisfying $(r_{\sigma(1)}^{\prime},\ldots,r_{\sigma(s)}^{\prime})=(r_{11},r_{12},\ldots,r_{mk_{m}})$. We call $l(G)$ the subdegree sequence of $G$ (of length $s$) and the $s$-tuple $(H_{11},H_{12},\ldots,H_{m(k_{m}-1)},H_{mk_{m}})$ a primitive-constituent sequence of $G$. The $m$-tuple of integers $(k_{1},\dots,k_{m})$ is called an intrinsic multiplicity sequence of $G$.
3.2. Abelian subgroups and canonical bounds
The following bound for abelian subgroups of the automorphism groups of smooth forms plays a key role in our study.
Theorem 3.4.
Let $F=F(x_{1},\dots,x_{r})$ be a smooth homogeneous polynomial of degree $d\geq 3$, where $r\geq 2$. Fix positive integers $r_{1},\dots,r_{m}$, where $1\leq m\leq r$. Let $N$ be an abelian subgroup of ${\rm Aut}(F)$. Suppose $N$ is generated by matrices of the form ${\rm diag}(\lambda_{1}I_{r_{1}},\lambda_{2}I_{r_{2}},\dots,\lambda_{m}I_{r_{m}})$. Then there exists an integer $k\geq 1$ and a partition $m=b+\sum_{i=1}^{k}a_{i}$ for some integers $b\geq 0$, $a_{i}>0$ ($i=1,\ldots,k$) such that $|N|$ divides $(d-1)^{b}\prod_{i=1}^{k}((d-1)^{a_{i}}-(-1)^{a_{i}})$. In particular, $|N|\leq d^{m}$.
Proof.
We set $s_{0}:=0$ and $s_{i}:=\sum_{j=1}^{i}r_{j}$ for $i=1,\dots,m$. Let $M=\{1,2,\dots,m\}$ and
$\mathcal{P}:=\{(i,j)\in M\times M\mid\;x_{s_{i}}^{d-1}x_{k_{j}}\in F\text{ forsome }k_{j}\;\text{ satisfying }s_{j-1}<k_{j}\leq s_{j}\}.$
By smoothness of $F$ and [OY19, Proposition 3.3], the projection map $p_{1}:\mathcal{P}\rightarrow M$ given by $(i,j)\mapsto i$ must be surjective. Consider the subgroup
$G_{\mathcal{P}}^{(m)}:=\{(\lambda_{1},\lambda_{2},\dots,\lambda_{m})|\;\lambda_{i}^{d-1}\lambda_{j}=1\text{ if }(i,j)\in\mathcal{P}\}\subset{{\mathbb{C}}^{\times}}\times\cdots\times{\mathbb{C}}^{\times}.$
Let $A={\rm diag}(\lambda_{1}I_{r_{1}},\lambda_{2}I_{r_{2}},\dots,\lambda_{m}I_{r_{m}})\in N$. Since $N$ is a subgroup of ${\rm Aut}(F)$, we have $A(F)=F$, which implies that $\lambda_{i}^{d-1}\lambda_{j}=1$ if $x_{s_{i}}^{d-1}x_{k_{j}}\in F$ and $s_{j-1}<k_{j}\leq s_{j}$. Then $N$ is isomorphic to a subgroup of $G_{\mathcal{P}}^{(m)}$ by the well-defined injective morphism $N\rightarrow G_{\mathcal{P}}^{(m)}$, $A\mapsto(\lambda_{1},\dots,\lambda_{m})$.
We may choose a subset $\mathcal{P}^{\prime}$ of $\mathcal{P}$ such that the cardinality of $\mathcal{P}^{\prime}$ is $m$ and $p_{1}(\mathcal{P}^{\prime})=M$. Note that $G_{\mathcal{P}}^{(m)}$ is a subgroup of $G_{\mathcal{P}^{\prime}}^{(m)}$. We claim that there exists an integer $k\geq 1$ and a partition $m=b+\sum_{i=1}^{k}a_{i}$ for some integers $b\geq 0$, $a_{i}>0$ ($i=1,\ldots,k$) such that
$|G_{\mathcal{P}^{\prime}}^{(m)}|=(d-1)^{b}\prod_{i=1}^{k}((d-1)^{a_{i}}-(-1)^{a_{i}}).$
Since $|N|$ divides $|G_{\mathcal{P}^{\prime}}^{(m)}|$ and the value on the right hand side is clearly at most $d^{m}$, it suffices to prove the claim. Let $p_{2}:\mathcal{P}^{\prime}\rightarrow M$ be the projection map given by $(i,j)\mapsto j$. There are two cases: (1) $p_{2}$ is surjective; (2) $p_{2}$ is not surjective.
For case (1), let $k$ be the largest integer such that $M$ can be expressed as a disjoint union of $k$ subsets $M_{i}$ ($i=1\dots,k$) with
$\mathcal{P}^{\prime}=\bigcup_{i=1}^{k}\mathcal{P}^{\prime}\cap(M_{i}\times M_{i}).$
We set $\mathcal{P}^{\prime}_{i}:=\mathcal{P}^{\prime}\cap(M_{i}\times M_{i})$. Then the cardinalities of $\mathcal{P}^{\prime}_{i}$ and $M_{i}$ are equal and we have $G_{\mathcal{P}^{\prime}}^{(m)}\cong G_{\mathcal{P}^{\prime}_{1}}^{(a_{1})}\times\cdots\times G_{\mathcal{P}^{\prime}_{k}}^{(a_{k})}$, where $a_{i}$ is the cardinality of the set $M_{i}$. Note that $m=\sum_{i=1}^{k}a_{i}$. By direct calculation, $|G_{\mathcal{P}^{\prime}_{i}}^{(a_{i})}|=(d-1)^{a_{i}}-(-1)^{a_{i}}$ and $|G_{\mathcal{P}^{\prime}}^{(m)}|=\prod_{i=1}^{k}((d-1)^{a_{i}}-(-1)^{a_{i}})$.
For case (2), we may assume $m\notin p_{2}(\mathcal{P}^{\prime})$. If $m=2$, then $G_{\mathcal{P}^{\prime}}^{(2)}=\{(\lambda_{1},\lambda_{2})|\;\lambda_{1}^{d}=\lambda_{2}^{d-1}\lambda_{1}=1\}$ of order $d(d-1)$. To prove the case $m\geq 3$ inductively, we set $\mathcal{P}^{\prime\prime}:=\mathcal{P}^{\prime}\cap(M^{\prime}\times M^{\prime})$, where $M^{\prime}=M\setminus\{m\}$. Then we have
$\mathcal{P}^{\prime}=\{(m,j_{m})\}\cup\mathcal{P}^{\prime\prime},$
where $j_{m}\in M^{\prime}$. The claim holds for $G_{\mathcal{P}^{\prime\prime}}^{(m-1)}$ by either case (1) or the induction hypothesis. Clearly, there is a short exact sequence
$1\rightarrow C_{d-1}\rightarrow G_{\mathcal{P}^{\prime}}^{(m)}\rightarrow G_{\mathcal{P}^{\prime\prime}}^{(m-1)}\rightarrow 1,$
where the third morphism is given by $(\lambda_{1},\dots,\lambda_{m})\mapsto(\lambda_{1},\dots,\lambda_{m-1})$. Thus $|G_{\mathcal{P}^{\prime}}^{(m)}|=(d-1)\cdot|G_{\mathcal{P}^{\prime\prime}}^{(m-1)}|$, which implies the claim for $G_{\mathcal{P}^{\prime}}^{(m)}$ since the claim holds for $G_{\mathcal{P}^{\prime\prime}}^{(m-1)}$. This completes the proof of the claim and hence the theorem. ∎
Now we are ready to introduce a bound for groups of automorphisms of smooth forms.
Definition 3.5.
Let $G$ be a subgroup of the automorphism group ${\rm Aut}(F)$ of a smooth form $F=F(x_{1},\ldots,x_{r})$ of degree $d\geq 3$, where $r\geq 1$. Let $(H_{1},\dots,H_{s})$ be a primitive-constituent sequence of $G$ with an intrinsic multiplicity sequence $(k_{1},\dots,k_{m})$. We define the canonical bound of $G$ by
$B(G,F)=d^{s}\prod_{i=1}^{s}|H_{i}|\prod_{j=1}^{m}{k_{j}!}.$
To simplify notation, we often use $B(G)$ to denote $B(G,F)$ if there is no confusion.
Combining Theorem 3.4 with the two associated exact sequences (3.1) and (3.2) of finite linear groups, we have the following observation.
Lemma 3.6.
Let $F=F(x_{1},\ldots,x_{r})$ be a smooth form of degree $d\geq 3$, where $r\geq 1$. If $G$ is a subgroup of ${\rm Aut}(F)$, then $|G|\leq B(G,F)$.
The following example gives an illustration for computing canonical bounds.
Example 3.7.
Following [OY19, Example 2.1 (17)], let $F=((x_{1}^{4}+x_{2}^{4})+(2+4\xi_{3}^{2})x_{1}^{2}x_{2}^{2})x_{3}+(-(x_{1}^{4}+x_{2}^{4})+(2+4\xi_{3}^{2})x_{1}^{2}x_{2}^{2})x_{4}+x_{3}^{4}x_{4}+x_{4}^{4}x_{3}+x_{5}^{5}$ and the subgroup $\widetilde{G}$ of ${\rm Aut}(F)$
$\widetilde{G}=\left\langle\begin{pmatrix}\xi_{8}^{3}&0&0&0&0\\ 0&\xi_{8}&0&0&0\\ 0&0&0&1&0\\ 0&0&1&0&0\\ 0&0&0&0&1\\ \end{pmatrix},\begin{pmatrix}-\frac{1}{\sqrt{2}}\xi_{8}&\frac{1}{\sqrt{2}}\xi_{8}&0&0&0\\ \frac{1}{\sqrt{2}}\xi_{8}^{3}&\frac{1}{\sqrt{2}}\xi_{8}^{3}&0&0&0\\ 0&0&\xi_{3}&0&0\\ 0&0&0&\xi_{3}^{2}&0\\ 0&0&0&0&1\\ \end{pmatrix},\begin{pmatrix}-1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\\ 0&0&0&1&0\\ 0&0&0&0&\xi_{5}\\ \end{pmatrix}\right\rangle.$
We denote by $(e_{1},e_{2},e_{3},e_{4},e_{5})$ the underlying basis of ${\mathbb{C}}^{5}$. Then using notations in $\bf{Set\textrm{-}up}$ 3.1, we have ${\mathbb{C}}^{5}=V_{1}\oplus V_{2}\oplus V_{3}$, $V_{1}=W_{11}$, $V_{2}=W_{21}\oplus W_{22}$, $V_{3}=W_{31}$, where $V_{1}=\langle e_{1},e_{2}\rangle$, $V_{2}=\langle e_{3},e_{4}\rangle$, $V_{3}=\langle e_{5}\rangle$, $W_{21}=\langle e_{3}\rangle$, $W_{22}=\langle e_{4}\rangle$. The linear group $\widetilde{G}$ has $4$ primitive constituents $H_{ij}$. Moreover, $S_{4}\cong H_{11}\subset{\rm PGL}(2,{\mathbb{C}})$, $|H_{21}|=|H_{22}|=|H_{31}|=1$, the subdegree sequence $l(\widetilde{G})=(2,1,1,1)$, and $\widetilde{G}$ has an intrinsic multiplicity sequence $(1,2,1)$. Then we have $B(\widetilde{G},F)=5^{4}\cdot 24\cdot 2=30000>480=|\widetilde{G}|$, as predicted by Lemma 3.6.
The following simple fact is useful for our study.
Lemma 3.8.
Let $k_{1},\ldots,k_{m}$ be positive integers with $m\geq 1$. Then $\prod_{j=1}^{m}{k_{j}!}\leq(k_{1}+\cdots+k_{m})!.$ The equality occurs if and only if $m=1$.
We say an invertible matrix $A$ is semi-permutation if $A$ is a diagonal matrix up to permutation of columns. Note that up to linear change of coordinates, a finite linear group consists of semi-permutation matrices if and only if its subdegree sequence is $(1,\ldots,1)$. As an immediate consequence of Lemma 3.6, we obtain an optimal bound for semi-permutation automorphism groups of smooth hypersurfaces.
Theorem 3.9.
Let $F=F(x_{1},\dots,x_{r})$ be a smooth form of degree $d\geq 3$ and $r\geq 2$. If ${\rm Aut}(F)$ consists of semi-permutations, then $|{\rm Aut}(F)|\leq d^{r}\cdot r!$. The equality occurs if and only if $F$ is isomorphic to the Fermat form $F_{d}^{r-2}=x_{1}^{d}+x_{2}^{d}+\cdots+x_{r}^{d}$.
Proof.
Since ${\rm Aut}(F)$ has the subdegree sequence $(1,\ldots,1)$ of length $s=r$ and the intrinsic multiplicity sequence $(k_{1},\ldots,k_{m})$ with $k_{1}+\cdots+k_{m}=r$, by Lemmas 3.6 and 3.8, we have
$|{\rm Aut}(F)|\leq B({\rm Aut}(F),F)=d^{r}\prod_{j=1}^{m}{k_{j}!}\leq d^{r}\cdot r!.$
If $|{\rm Aut}(F)|=d^{r}\cdot r!$, then $N\cong C_{d}^{r}$, $m=1$, and $K_{1}\cong S_{r}$ (here $N$ and $K_{1}$ are as in $\mathbf{Set{\textrm{-}}up}$ 3.1 and Definition 3.2). In this case, $\{x_{i}^{d}\}_{1\leq i\leq r}$ are the only monomials preserved by $N$, thus $F$ is a Fermat form, up to replacing $x_{i}$ by their non-zero multiples. ∎
4. Fermat-test ratios and proof of most cases
A famous theorem of Jordan states that, for every integer $r\geq 1$, there exists a constant such that any finite subgroup of ${\rm GL}(r,{\mathbb{C}})$ has an abelian normal subgroup of index at most this constant. Consequently, for each integer $r\geq 1$, there are only finitely many finite primitive groups in ${\rm PGL}(r,{\mathbb{C}})$ (up to conjugation), and Collins [Col08a, Theorem A] determined the maximum order ${\rm JC}(r)$ among such groups. Based on canonical bounds and the values of ${\rm JC}(r)$, we introduce the notion of Fermat-test ratio (Definition 4.5) and we show that Fermat-test ratios have many nice properties which will be frequently used in our study. Using Fermat-test ratios, we quickly prove that our main result Theorem 1.1 holds if $n\geq 26$ or $d\geq 18$ (Theorem 4.13).
As a generalization of subdegree sequences of linear groups, the following definition will be used in the definiton of Fermat-test ratios.
Definition 4.1.
[…]
Remark 4.2.
We allow $r^{0}$ to appear in the exponential type of $l$ even if $r$ is not a subdegree of $l$. Under such convention, the equations in Definition 4.1 still hold. For example, by $l$ being the subdegree sequence of the exponential type $5^{0}3^{2}2^{0}1^{3}$ we mean $l=(3,3,1,1,1)$.
For later use, we define sum/difference between subdegree sequences.
Definition 4.3.
Let $l_{1}$ and $l_{2}$ be two subdegree sequences of exponential types $r_{1}^{a_{1}}r_{2}^{a_{2}}\cdots r_{m}^{a_{m}}$ and $r_{1}^{b_{1}}r_{2}^{b_{2}}\cdots r_{m}^{b_{m}}$, where $a_{i},b_{i}\geq 0$, $m\geq 1$, $r_{1}>\cdots>r_{m}\geq 1$. We denote by $l_{1}+l_{2}$ the subdegree sequence of the exponential type $r_{1}^{a_{1}+b_{1}}r_{2}^{a_{2}+b_{2}}\cdots r_{m}^{a_{m}+b_{m}}$. If $a_{i}\geq b_{i}$ for all $i$, we denote by $l_{1}-l_{2}$ the subdegree sequence of the exponential type $r_{1}^{a_{1}-b_{1}}r_{2}^{a_{2}-b_{2}}\cdots r_{m}^{a_{m}-b_{m}}$.
Example 4.4.
Let $G$ be as in Example 3.7. Then the subdegree sequence $l(G)$ has the exponential type $2^{1}1^{3}$, and the multiplicity sequence of $l(G)$ is $(1,3)$ which is different from the intrinsic multiplicity sequence $(1,2,1)$ of $G$.
We recall the values of the function ${\rm JC}(r)$ ([Col08a, Theorem A]). For $r=10,11$ or $r\geq 13$, ${\rm JC}(r)=(r+1)!$; for the remaining cases, ${\rm JC}(r)$ is as follows:
$r$ 1 2 3 4 5 6 7 8 9 12
${\rm JC}(r)$ 1 60 360 25920 25920 6531840 1451520 348364800 4199040 448345497600
Now we give the definition of Fermat-test ratio which is crucial for our classification of smooth hypersurfaces with large automorphism groups.
Definition 4.5.
Let $l=(r_{1}^{\prime},\ldots,r_{s}^{\prime})$ be a subdegree sequence of length $s\geq 1$ with the exponential type $r_{1}^{k_{1}}r_{2}^{k_{2}}\cdots r_{m}^{k_{m}}$ and total degree $r\geq 1$. For an integer $d\geq 3$, we define the Fermat-test ratio of the pair $(l,d)$ by
$R(l,d)=\frac{d^{s}\cdot\prod\limits_{i=1}^{s}{{\rm JC}(r_{i}^{\prime})}\cdot\prod\limits_{j=1}^{m}{k_{j}!}}{d^{r}\cdot r!}.$
Let $\mathcal{H}=(H_{1},\ldots,H_{s})$ be an $s$-tuple of finite primitive subgroups $H_{i}\subset{\rm PGL}(r_{i}^{\prime},{\mathbb{C}})$ with the subdegree sequence $l(\mathcal{H})=l$. We define the Fermat-test ratio of the pair $(\mathcal{H},d)$ by
$R(\mathcal{H},d)=\frac{d^{s}\cdot\prod\limits_{i=1}^{s}{|H_{i}|}\cdot\prod\limits_{j=1}^{m}{k_{j}!}}{d^{r}\cdot r!}.$
Let $F=F(x_{1},\ldots,x_{r})$ be a smooth form of degree $d\geq 3$, where $r\geq 1$. We call $R(l({\rm Aut}(F)),d)$ the Fermat-test ratio of $F$, denoted by $R(F)$.
By Lemmas 3.6, 3.8 and definition of Fermat-test ratios, we have the following observation.
Lemma 4.6.
Let $F=F(x_{1},\ldots,x_{r})$ be a smooth form of degree $d\geq 3$ and $r\geq 1$. Let $\mathcal{H}:=(H_{1},\ldots,H_{s})$ be a primitive-constituent sequence of ${\rm Aut}(F)$ with $r_{1}\geq\cdots\geq r_{m}$, where $r_{i}$ is the degree of $H_{i}$. Then we have
$R(F)\geq\frac{R(F)\cdot|H_{i}|}{{\rm JC}(r_{i})}\geq R(\mathcal{H},d)\geq\frac{B({\rm Aut}(F))}{d^{r}\cdot r!}\geq\frac{|{\rm Aut}(F)|}{d^{r}\cdot r!}.$
In particular, if $|{\rm Aut}(F)|\geq d^{r}\cdot r!$, then $R(F)\geq R(\mathcal{H},d)\geq 1$.
By direct calculation, we find some basic properties of Fermat-test ratios.
Lemma 4.7.
Let $l$ be a subdegree sequence of length $s\geq 1$. Let $d,d^{\prime}$ be integers at least $3$. Then $R(l,d)/R(l,d^{\prime})=(d^{\prime}/d)^{v(l)-s}$. In particular, if $d>d^{\prime}$, then $R(l,d)\leq R(l,d^{\prime})$.
The properties of Fermat-test ratios in Lemmas 4.8, 4.9 and 4.11 make computing/bounding values of Fermat-test ratios quite feasible.
Lemma 4.8.
Let $l_{1}$, $l_{2}$ be subdegree sequences and let $d\geq 3$ be a positive integer. Then we have
  1. (1)
    $\tbinom{v({l_{1}+l_{2}})}{v({l_{1}})}\cdot R(l_{1}+l_{2},d)\geq R(l_{1},d)\cdot R(l_{2},d)$. The equality occurs if and only if the subdegrees of $l_{1}$ and $l_{2}$ are disjoint;
  2. (2)
    Suppose the subdegrees of $l_{1}$ and $l_{2}$ are disjoint, if $v(l_{2})\geq v(l_{2}^{\prime})$ and $R(l_{2},d)\leq R(l_{2}^{\prime},d)$, then $R(l_{1}+l_{2},d)\leq R(l_{1}+l_{2}^{\prime},d)$; and
  3. (3)
    Let $s_{i}$ be the length of $l_{i}$ ($i=1,2$). Suppose $s_{1}\geq s_{2}$ and $v(l_{1})\leq v(l_{2})$. If $R(l_{1},d)>R(l_{2},d)$, then $R(l_{1},d+1)>R(l_{2},d+1)$.
Proof.
The three statements are following from direct calculations. ∎
Lemma 4.9.
Let $l$ be a subdegree sequence containing $r_{0}>1$ with multiplicity $k_{0}\geq 1$. Then, for $d\geq 3$, $R(l,d)>R((l+(r_{0}),d))$ holds if one of the following conditions is true:
  1. (1)
    $r_{0}=2$ and $k_{0}\geq 5$;
  2. (2)
    $r_{0}=4$ and $k_{0}\geq 2$;
  3. (3)
    $r_{0}=3$ or $r_{0}\geq 5$.
In particular, if $l$ is of exponential type $r_{0}^{k_{0}}$ with $r_{0}\notin\{2,3,4,5,6,8\}$, then $R(l,d)<1$.
Proof.
Let $l_{2}$ be the subdegree sequence of the exponential type $r_{0}^{k_{0}}$. We set $l_{1}:=l-l_{2}$. By (1) in Lemma 4.8 we have
$\tbinom{v({l})+r_{0}}{(k_{0}+1)r_{0}}R(l+(r_{0}),d)=R(l_{1},d)R(l_{2}+(r_{0}),d),\;\;\tbinom{v({l})}{k_{0}r_{0}}R(l,d)=R(l_{1},d)R(l_{2},d).$
Then the ratio
$\lambda:=\frac{R(l+(r_{0}),d)}{R(l,d)}=\frac{\tbinom{v({l})}{k_{0}r_{0}}}{\tbinom{v({l})+r_{0}}{(k_{0}+1)r_{0}}}\cdot\frac{R(l_{2}+(r_{0}),d)}{R(l_{2},d)}=\frac{v(l)!}{(v(l)+r_{0})!}\cdot\frac{{\rm JC}(r_{0})(k_{0}+1)}{d^{r_{0}-1}}.$ (4.1)
Since $\lambda$ decreases as $v(l_{1})$ increases or $d$ increases, we may take $v(l_{1})=0$ and $d=3$. Then
$\lambda\leq\frac{(r_{0}k_{0})!}{(r_{0}(k_{0}+1))!}\cdot\frac{(k_{0}+1){\rm JC}(r_{0})}{3^{r_{0}-1}}=\frac{(r_{0}k_{0})!}{(r_{0}(k_{0}+1)-1)!}\cdot\frac{{\rm JC}(r_{0})}{r_{0}3^{r_{0}-1}}.$
We notice that the value in the right hand side of the inequality decreases as $k_{0}$ increases. By direct calculation, if $r_{0}=2$, $k_{0}=5$, we have $\lambda\leq\frac{10!\cdot 60}{11!\cdot 2\cdot 3}=\frac{10}{11}<1$. Thus (1) follows. Similarly, one may check if $(r_{0},k_{0})=(3,1)$, $(4,2)$, $(5,1)$, $(6,1)$, $(7,1)$, $(8,1)$, $(9,1)$, $(12,1)$, we have $\lambda<1$.
For $r_{0}=10,11$ or $r_{0}\geq 13$, we have $\lambda<\frac{(r_{0}+1)!}{3^{r_{0}-1}r_{0}^{r_{0}}}<\frac{r_{0}+1}{3^{r_{0}-1}}\cdot\frac{r_{0}^{r_{0}}}{r_{0}^{r_{0}}}<1$. By direct calculation, we find that $R((r_{0}),d)\leq R((r_{0}),3)<1$ for $r_{0}=7$, or $r_{0}\geq 9$. Therefore, $R(r_{0}^{k_{0}},d)<1$ if $r_{0}=7$ or $r_{0}\geq 9$. ∎
Remark 4.10.
$R(3^{1},3)=20/3$, $R(4^{1},3)=40$, $R(4^{2},3)=320/7$, $R(5^{1},3)=8/3$, $R(6^{1},3)=112/3$, $R(8^{1},3)=320/81$.
Lemma 4.11.
Let $d\geq 3$ and $m\geq 1$ be integers. Let $l_{1}$ be a subdegree sequence of exponential type $r_{1}^{k_{1}}\cdots r_{m}^{k_{m}}$ with $r_{i}>1$, $k_{i}>0$ and $R(r_{i}^{k_{i}},d)\geq 1$ for all $1\leq i\leq m$. Then we have $R(l_{1},d)\leq R(l_{1}^{\prime},d)$, where
$l_{1}^{\prime}\text{ is of exponential type}\begin{cases}4^{1}&\quad\text{if }v(l_{1})=4~{}~{}and~{}~{}d=3\\ 2^{a}{~{}~{}with~{}~{}}a=v(l_{1})/2&\quad\text{if }v(l_{1}){~{}~{}is~{}~{}even~{}~{}and~{}~{}}(v(l_{1}),d)\neq(4,3)\\ 2^{a}{~{}~{}with~{}~{}}a=(v(l_{1})-1)/2&\quad\text{if }v(l_{1}){~{}~{}is~{}~{}odd}\\ \end{cases}$
In particular, if $v(l_{1})\geq 28$, we have $R(l_{1},d)<1$.
Proof.
By Lemma 4.9, we have $r_{i}\in\{2,3,4,5,6,8\}$ for all $i$. Then we may assume $l_{1}$ is of exponential type $8^{a_{8}}6^{a_{6}}5^{a_{5}}4^{a_{4}}3^{a_{3}}2^{a_{2}}$ with $a_{i}\geq 0$. Furthermore, as $R(2^{14},3)$, $R(3^{4},3)$, $R(4^{4},3)$, $R(5^{2},3)$, $R(6^{3},3)$, and $R(8^{2},3)$ are all less than 1, applying Lemmas 4.7 and 4.9, we have $a_{8}\leq 1$, $a_{6}\leq 2$, $a_{5}\leq 1$, $a_{4}\leq 3$, $a_{3}\leq 3$, $a_{2}\leq 13$. By calculation, we observe $R(5^{1}3^{1},3)<R(2^{4},3)$, $R(5^{1}3^{3},3)<R(2^{7},3)$, $R(3^{1},3)<R(2^{1},3)$, $R(3^{2},3)<R(2^{3},3)$, $R(3^{3},3)<R(2^{4},3)$, $R(4^{2},3)<R(2^{4},3)$, $R(4^{3},3)<R(2^{6},3)$, $R(5^{1},3)<R(2^{2},3)$, $R(6^{1},3)<R(2^{3},3)$, $R(6^{2},3)<R(2^{6},3)$ and $R(8^{1},3)<R(2^{4},3)$.
If $a_{4}\neq 1$, then by Lemma 4.8 (2)-(3) and replacing subdegrees $3,4,5,6,8$ by $2$ with suitable multiplicities, we have $R(l_{1},d)\leq R(2^{a_{2}^{\prime}},d)$ with $a_{2}^{\prime}=\frac{v(l_{1})}{2}$ (resp. $a_{2}^{\prime}=\frac{v(l_{1})-1}{2}$) when $v(l_{1})$ is even (resp. odd).
If $a_{4}=1$, then similar to the previous case, we have $R(l_{1},d)\leq R(4^{1}2^{b_{2}},d)$ with $b_{2}=\frac{v(l_{1})-4}{2}$ (resp. $b_{2}=\frac{v(l_{1})-5}{2}$) when $v(l_{1})$ is even (resp. odd). Since the ratio $R(4^{1}2^{b_{2}},d)/R(2^{b_{2}+2},d)$ $=\frac{25920}{3600d(b_{2}+1)(b_{2}+2)}$, we find that $R(4^{1}2^{b_{2}},d)\leq R(2^{b_{2}+2},d)$ unless $(b_{2},d)=(0,3)$. Thus, we have $R(l_{1},d)\leq R(2^{b_{2}+2},d)$ unless $(v(l_{1}),d)=(4,3)$.
Moreover, since $R(2^{14},d)<1$ for $d=3$, by Lemmas 4.7 and 4.9, we have $R(l_{1},d)<1$ for all $l_{1}$ satisfying $v(l_{1})\geq 28$. ∎
Remark 4.12.
For later use, here we present (in decreasing order) the integer parts $[R(l_{1}^{\prime},3)]$ of the values $R(l_{1}^{\prime},3)$ for all possible $l_{1}^{\prime}$ with $v(l_{1}^{\prime})\leq 26$ in Lemma 4.11: $[R(2^{5},3)]=105$, $[R(2^{6},3)]=96$, $[R(2^{4},3)]=95$, $[R(2^{7},3)]=74$, $[R(2^{3},3)]=66$, $[R(2^{8},3)]=49$, $R(4^{1},3)=40$, $[R(2^{2},3)]=33$, $[R(2^{9},3)]=29$, $[R(2^{10},3)]=15$, $R(2^{1},3)=10$, $[R(2^{11},3)]=7$, $[R(2^{12},3)]=3$, $[R(2^{13},3)]=1$.
Next we use Fermat-test ratios to prove that our main result Theorem 1.1 holds for all possible pairs $(n,d)$ except only finitely many cases.
Theorem 4.13.
Let $F=F(x_{1},\dots,x_{n+2})$ be a smooth form of degree $d$, where $n\geq 1$, $d\geq 3$. Suppose $X_{F}$ is not isomorphic to the Fermat hypersurface in ${\mathbb{P}}^{n+1}$ of degree $d$. If $n\geq 26$ or $d\geq 18$, then $|{\rm Aut}(X_{F})|<d^{n+1}\cdot(n+2)!$.
Proof.
Suppose the subdegree sequence $l:=l({\rm Aut}(F))$ of ${\rm Aut}(F)$ is of exponential type $r_{1}^{k_{1}}\cdots r_{m}^{k_{m}}$ with $m\geq 1$ and $k_{i}>0$ for all $i$. By Theorem 3.9, we may assume $r_{1}>1$. Then by Lemma 4.6, it suffices to show $R(l,d)<1$. Let $\mathcal{A}=\{i|\;r_{i}>1,R(r_{i}^{k_{i}},d)\geq 1,1\leq i\leq m\}.$ We define $l_{1}$ to be the sum of all the subdegree sequences of exponential types $r_{j}^{k_{j}}$ ($j\in\mathcal{A}$). Let $l_{2}:=l-l_{1}$.
Case $n\geq 26$. If $l=l_{1}$ or $l=l_{2}$, then by Lemmas 4.11 and 4.8, $R(l,d)<1$. Then we may assume that $l\neq l_{1},l_{2}$ and $v(l_{1})<28$. Since $n+2=v(l_{1}+l_{2})\geq 28$, by Lemmas 4.8, 4.11 and Remark 4.12, we have
$R(l_{1}+l_{2},d)=\frac{R(l_{1},d)\cdot R(l_{2},d)}{\tbinom{v({l_{1}+l_{2}})}{v({l_{1}})}}\leq\frac{R(l_{1}^{\prime},d)}{\tbinom{v({l_{1}+l_{2}})}{v({l_{1}})}}\leq\frac{R(l_{1}^{\prime},d)}{\tbinom{28}{v({l_{1}})}}\leq\frac{R(l_{1}^{\prime},3)}{\tbinom{28}{v({l_{1}})}}<1.$
Case $d\geq 18$. By Lemma 4.7, it suffices to show $R(l,18)<1$. For positive integers $r_{0}>1$ and $k_{0}>0$, by Lemma 4.9 and direct calculation, $R(r_{0}^{k_{0}},18)\geq 1$ if and only if $(r_{0},k_{0})=(2,1)$. Thus, by Lemma 4.11, either $l_{1}=l_{1}^{\prime}$ of exponential type $2^{1}$ or $l=l_{2}$, which implies $R(l,18)<1$ by Lemma 4.8 again. This completes the proof of the theorem. ∎
By similar arguments in the proof of Theorem 4.13, we obtain additional properties of Fermat-test ratios, which will be used in the subsequent sections.
Lemma 4.14.
Let $a\geq 0$ be the multiplicity of $1$ in a subdegree sequence $l$ and $d\geq 3$. Then
  1. (1)
    $R(l,d)<106$.
  2. (2)
    If $l$ has at least two distinct subdegrees, then $R(l,d)<60$.
  3. (3)
    If $a\geq 2$ (resp. $a\geq 1$), then $R(l,d)<3$ (resp. $R(l,d)<11$).
Remark 4.15.
Let $d\geq 3$, $a\geq 2$. Let $l$ be a subdegree sequence containing 1 and $a$ with $R:=R(l,d)\geq 1$. Then $a\in\{2,3,4,6\}$. Moreover, if $a=3$ (resp. $6$), then $(l,d,R)=(3^{1}2^{1}1^{1},3,10/9)$, $(3^{1}1^{1},3,5/3)$ (resp. $(6^{1}1^{1},4,81/64)$, $(6^{1}1^{2},3,4/3)$, $(6^{1}2^{1}1^{1},3,40/27)$, $(6^{1}1^{1},3,16/3)$).
5. Refined bounds via special monomials
In this section, based on canonical bounds, we derive several refined upper bounds for the size of automorphism groups of smooth forms containing monomials involving different primitive constituents (Lemmas 5.4, 5.5, 5.6, 5.7). Such bounds will be frequently used in Section 7.
Smoothness of a polynomial $F$ sometimes implies existence of monomials of special shape.
Lemma 5.1.
Let $F=F(x_{1},\ldots,x_{k+a})$ be a smooth form of degree $d\geq 3$ with $k\geq 2$ and $a>0$. We write $F=F_{1}+F_{2}$, where $F_{1}\in{\mathbb{C}}[x_{1},\ldots,x_{k}]$ and $F_{2}\in(x_{k+1},\ldots,x_{k+a})\cdot{\mathbb{C}}[x_{1},\ldots,x_{k+a}]$ are forms of degree $d$. Suppose $F_{1}$ is not smooth. Then $F_{2}\notin(x_{k+1},\ldots,x_{k+a})^{2}\cdot{\mathbb{C}}[x_{1},\ldots,x_{k+a}]$. In particular, there exist integers $d_{1},\ldots,d_{k}\geq 0$ and a monomial $\mathfrak{m}\in F$ with $\mathfrak{m}=x_{1}^{d_{1}}\cdots x_{k}^{d_{k}}\cdot x_{k+j}$ for some $1\leq j\leq a$.
Proof.
Since $F_{1}$ is not smooth, we may assume all partial derivatives of $F_{1}$ vanishes at $(1,0,\dots,0)$ $\in{\mathbb{C}}^{k}$. Then by direct computation, if $F_{2}\in(x_{k+1},\ldots,x_{k+a})^{2}\cdot{\mathbb{C}}[x_{1},\ldots,x_{k+a}]$, all partial derivatives of $F$ vanishes at $(1,0,\ldots,0)\in{\mathbb{C}}^{k+a}$, a contradiction. Therefore, $F_{2}\notin(x_{k+1},\ldots,x_{k+a})^{2}\cdot{\mathbb{C}}[x_{1},\ldots,x_{k+a}]$, which implies the last statement in the lemma. ∎
The following lemma will be used in the proof of Lemmas 5.4 and 5.5.
Lemma 5.2.
Let $F=F(x_{1},\dots,x_{n+2})$ be a form of degree $d\geq 3$ with $n\geq 1$ and $G$ be a finite subgroup of ${\rm Aut}(F)$. Suppose the underlying basis of ${\mathbb{C}}^{n+2}$ is compatible with a decomposition ${\mathbb{C}}^{n+2}=V_{1}\oplus V_{2}$, where $V_{i}$ are $G$-stable subspaces of dimension $n_{i}\geq 1$ ($i=1,2$). If the $G$-space $V_{2}$ is irreducible and the $S^{d-1}(V_{1}^{*})\otimes S^{1}(V_{2}^{*})$-component $F^{(d-1,1)}$ of $F$ is not zero, then we have
$F^{(d-1,1)}=F^{\prime}_{1}x_{n_{1}+1}+\cdots+F^{\prime}_{n_{2}}x_{n_{1}+n_{2}},$
[…]
Proof.
Recall that $F^{(d-1,1)}\in S^{d-1}(V_{1}^{*})\otimes S^{1}(V_{2}^{*})$. Let $W_{2}^{\prime}$ be the minimal subvector space of $V_{2}^{*}$ such that $F^{(d-1,1)}\in S^{d-1}(V_{1}^{*})\otimes S^{1}(W_{2}^{\prime})$. By $F^{(d-1,1)}\neq 0$, we have $W_{2}^{\prime}\neq 0$. Since $G$ preserves $F$ and $V_{i}$ $(i=1,2)$ are $G$-stable, we have $G$ preserves $F^{(d-1,1)}$, which implies $W_{2}^{\prime}$ is $G$-stable. From the irreducibility of $V_{2}$, we get that $V_{2}^{*}$ is also irreducible under the (right) action by $G$. Then $W_{2}^{\prime}=V_{2}^{*}$, which implies a decomposition of $F^{(d-1,1)}$ as in the lemma. ∎
For our purposes and conventions, we recall some notations in $\bf{Set\textrm{-}up}$ 3.1.
Set-up 5.3.
Let $F=F(x_{1},\dots,x_{r})$ be a smooth form of degree $d\geq 3$ with $r\geq 3$. Let $G$ be a subgroup of ${\rm Aut}(F)$. Let $V_{i},W_{ij},K_{i},k_{i},r_{ij},H_{ij},m,P,\widetilde{H}_{ij},\phi_{ij},\pi_{ij},\phi,N$ be as in $\bf{Set\textrm{-}up}$ 3.1 and Definition 3.2. Let $\tilde{\bar{H}}_{ij}:=\phi_{ij}(P)$ and $\bar{H}_{ij}:=\pi_{ij}(\tilde{\bar{H}}_{ij})$. We assume that the underlying basis of ${\mathbb{C}}^{r}$ is compatible with the decomposition ${\mathbb{C}}^{r}=\oplus_{i,j}W_{ij}$. The decomposition
$S^{d}({{\mathbb{C}}^{r}}^{*})=\mathop{\vbox{\hbox{{$\bigoplus$}}}}_{\sum_{i,j}d_{ij}=d}S^{d_{11}}(W_{11}^{*})\otimes S^{d_{12}}(W_{12}^{*})\otimes\cdots\otimes S^{d_{mk_{m}}}(W_{mk_{m}}^{*})$
naturally induces a decomposition
$F=\sum_{\sum_{i,j}d_{ij}=d}F^{(d_{11},d_{12},\ldots,d_{mk_{m}})}\in S^{d}({\mathbb{C}}^{r*}).$
We use $F_{ij}$ to denote the component (possibly being zero) in $S^{d}(W_{ij}^{*})$.
Note that $\tilde{\bar{H}}_{ij}$ and $\bar{H}_{ij}$ are normal subgroups of $\widetilde{H}_{ij}$ and $H_{ij}$ respectively since $P$ is normal in $G$. Clearly primitive constituents $H_{ij}$ of $G$ preserve $F_{ij}$. In the following lemmas and their proofs, we adopt the notations in $\bf{Set\textrm{-}up}$ 5.3.
Lemma 5.4.
Suppose $m\geq 2$ and let $V_{1}^{\prime}:=\oplus_{i\neq 2}V_{i}$. If there is a monomial $\mathfrak{m}\in S^{d-1}(V_{1}^{\prime*})\otimes S^{1}(V_{2}^{*})$ in $F$, then
$|G|\leq\frac{B(G,F)}{|H_{21}|^{k_{2}}}\leq\frac{R(l(G),d)\cdot d^{r}\cdot r!}{{\rm JC}(r_{21})^{k_{2}}}.$
Proof.
We have the following commutative diagram
[…]
Note that the three injections are inclusions of normal subgroups. Consider the exact sequence
$1\longrightarrow\widetilde{N}\longrightarrow P\longrightarrow H_{11}\times\cdots\times H_{1k_{1}}\times H_{31}\times\cdots\times H_{3k_{1}}\times\cdots\times H_{m1}\times\cdots\times H_{mk_{m}},$
where the last morphism is given by $g\mapsto([g|W_{11}],\ldots,[g|W_{1k_{1}}],[g|W_{31}],\ldots,[g|W_{mk_{m}}])$ and $\widetilde{N}$ is defined as its kernel. In order to prove the lemma, it suffices to show $\widetilde{N}\subset N$. Recall that the underlying basis $(e_{1},\ldots,e_{r})$ of ${\mathbb{C}}^{r}$ is compatible with the decomposition ${\mathbb{C}}^{r}=\oplus_{i,j}W_{ij}$. For any
$A={\rm diag}(\lambda_{11}I_{r_{11}},\dots,\lambda_{1k}I_{r_{1k_{1}}},A_{21},A_{22},\dots,A_{2k_{2}},\lambda_{31}I_{r_{31}},\dots,\lambda_{mk_{m}}I_{r_{mk_{m}}})\in\widetilde{N},$
$B={\rm diag}(B_{11},\dots,B_{1k_{1}},\dots,B_{m1},\dots,B_{mk_{m}})\in H,$
we have the commutator $C=[A,B]$ is the identity matrix, by the shape of $A$, $B$, $C(F)=F$ and Lemma 5.2.
Therefore, we have $\phi_{2j}(A)\in Z(\tilde{\bar{H}}_{2j})$ for any $1\leq j\leq k_{2}$. Since $\tilde{\bar{H}}_{2j}\vartriangleleft\widetilde{H}_{2j}$, we have $Z(\tilde{\bar{H}}_{2j})\vartriangleleft\widetilde{H}_{2j}$. Then by primitivity of $\widetilde{H}_{2j}$, we have
$Z(\tilde{\bar{H}}_{2j})\subset Z(\widetilde{H}_{2j})\subset\langle\{\lambda I_{r_{2j}}|\lambda\in{\mathbb{C}}^{\times}\}\rangle.$
Thus $A_{2j}\in\langle\{\lambda I_{r_{2j}}|\lambda\in{\mathbb{C}}^{\times}\}\rangle$ for any $1\leq j\leq k_{2}$. Therefore, $A\in N$ and the lemma follows. ∎
By similar arguments in Lemma 5.4, we obtain the following lemma.
Lemma 5.5.
Suppose $1\leq a\leq m$ and $k_{a}\geq 2$. If there is a monomial $\mathfrak{m}\in S^{d-1}(W_{a1}^{*})\otimes S^{1}(W_{a2}^{*})$ in $F$, then one of the following statements holds:
  1. (1)
    $k_{a}=2$ and $|G|\leq\frac{B(G,F)}{|H_{a1}|}\leq\frac{R(l(G),d)\cdot d^{r}\cdot r!}{{\rm JC}(r_{a1})}$;
  2. (2)
    $k_{a}\geq 3$ and $|G|\leq\frac{B(G,F)}{2\cdot|H_{a1}|}\leq\frac{R(l(G),d)\cdot d^{r}\cdot r!}{2\cdot{\rm JC}(r_{a1})}$.
Lemma 5.6.
Let $1\leq i_{1}\leq m$, $1\leq j_{1}\leq k_{i_{1}}$, $0<d_{1}<d$. Suppose there is a monomial $\mathfrak{m}\in S^{d_{1}}(W_{i_{1}j_{1}}^{*})\otimes S^{d-d_{1}}(\oplus_{(i_{2},j_{2})\neq(i_{1},j_{1})}W_{i_{2}j_{2}}^{*})$ in $F$. Then $\bar{H}_{i_{1}j_{1}}$ contains a normal subgroup $\bar{\bar{H}}_{i_{1}j_{1}}$ preserving a non-zero form of degree $d_{1}$ and
$|G|\leq\frac{B(G,F)}{[H_{i_{1}j_{1}}:\bar{\bar{H}}_{i_{1}j_{1}}]}\leq\frac{R(l(G),d)|\bar{\bar{H}}_{i_{1}j_{1}}|\cdot d^{r}\cdot r!}{{\rm JC}(r_{i_{1}j_{1}})}.$
Proof.
Without loss of generality, we assume $(i_{1},j_{1})=(1,1)$. Consider the exact sequence
$1\longrightarrow\widetilde{N}_{11}\longrightarrow P\longrightarrow H_{12}\times\cdots\times H_{1k_{1}}\times H_{21}\times\cdots\times H_{2k_{2}}\times\cdots\times H_{m1}\times\cdots\times H_{mk_{m}},$ (5.1)
where the last morphism is given by $g\mapsto([g|W_{12}],\ldots,[g|W_{1k_{1}}],[g|W_{21}],\ldots,[g|W_{mk_{m}}])$ and $\widetilde{N}_{11}$ is defined as its kernel. Let $\bar{\bar{H}}_{11}:=\phi_{11}(\widetilde{N}_{11})$. Then $\bar{\bar{H}}_{11}\vartriangleleft\bar{H}_{11}$ and we have the exact sequence
$1\longrightarrow N\longrightarrow\widetilde{N}_{11}\longrightarrow\bar{\bar{H}}_{11}\longrightarrow 1$ (5.2)
From the exact sequences (3.1), (3.2), (5.1), (5.2) and by Theorem 3.4, we have
$|G|\leq|P|\prod_{i}k_{i}!\leq\frac{|\widetilde{N}_{11}|\prod_{i}k_{i}!\prod_{i,j}|H_{ij}|}{|H_{11}|}\leq\frac{|N||\bar{\bar{H}}_{11}|\prod_{i}k_{i}!\prod_{i,j}|H_{ij}|}{|H_{11}|}\leq\frac{d^{s}\prod_{i}k_{i}!\prod_{i,j}|H_{ij}|}{[H_{11}:\bar{\bar{H}}_{11}]},$
which implies the inequalities in the lemma by definitions of canonical bounds and Fermat-test ratios. From the existence of $\mathfrak{m}$, we infer that $\bar{\bar{H}}_{11}$ preserves a non-zero form of degree $d_{1}$. ∎
Lemma 5.7.
Let $1\leq i_{1}\leq m$, $1\leq j_{1}\leq k_{i_{1}}$ and $r_{i_{1}j_{1}}\geq 2$. We define
$\mathcal{A}:=\{(i_{2},j_{2})|\;(i_{2},j_{2})\neq(i_{1},j_{1}),\;\exists\text{ a monomial }\mathfrak{m}\in S^{d-1}(W_{i_{1}j_{1}}^{*})\otimes S^{1}(W_{i_{2}j_{2}}^{*})\text{ in }F\}.$
Let $c$ be the cardinality of $\mathcal{A}$. If the form $F_{i_{1}j_{1}}$ is either zero or not smooth, then $c\neq 0$ and the following statements hold:
  1. (1)
    $|G|\leq\frac{B(G,F)}{|H_{i_{2}j_{2}}|}\leq\frac{R(l(G),d)\cdot d^{r}\cdot r!}{{\rm JC}(r_{i_{2}j_{2}})}$ for all $(i_{2},j_{2})\in\mathcal{A}$;
  2. (2)
    $|G|\leq\frac{B(G,F)}{d^{c-1}}\leq\frac{R(l(G),d)\cdot d^{r}\cdot r!}{d^{c-1}}$;
  3. (3)
    If $c=1$, $r_{i_{2}j_{2}}=1$ for $(i_{2},j_{2})\in\mathcal{A}$ and $G$ contains $\mu_{r,d}$, then $F_{i_{1}j_{1}}$ is not zero and $H_{i_{1}j_{1}}$ admits an $F_{i_{1}j_{1}}$-lifting preserving some non-zero form of degree $d-1$.
Proof.
We may assume $(i_{1},j_{1})=(1,1)$. We choose a basis of ${\mathbb{C}}^{r}$ compatible with the decomposition ${\mathbb{C}}^{r}=\oplus_{i,j}W_{ij}$. By the smoothness of $F$ and Lemma 5.1, we have $c\neq 0$. The statement (1) is a direct consequence of Lemmas 5.4 and 5.5.
For (2), let $A={\rm diag}(\lambda_{11}I_{r_{11}},\ldots,\lambda_{1k_{1}}I_{r_{1k_{1}}},\ldots,\lambda_{m1}I_{r_{m1}},\ldots,\lambda_{mk_{m}}I_{r_{mk_{m}}})\in N$. By $A(F)=F$, we get $\lambda_{11}^{d-1}\lambda_{i_{2}j_{2}}=\cdots=\lambda_{11}^{d-1}\lambda_{i_{c+1}j_{c+1}}=1$, where $\mathcal{A}=\{(i_{2},j_{2}),\ldots,(i_{c+1},j_{c+1})\}$. Thus, $\lambda_{i_{2}j_{2}}=\cdots=\lambda_{i_{c+1}j_{c+1}}$, which implies $|N|\leq d^{s-(c-1)}$ by Theorem 3.4. From this, we obtain the inequalities in (2) using the associated exact sequences (3.1) and (3.2).
For (3), we define $W^{\prime}:=\oplus_{(i^{\prime},j^{\prime})\neq(1,1),(i_{2},j_{2})}W_{i^{\prime}j^{\prime}}$, where $\mathcal{A}=\{(i_{2},j_{2})\}$. Let $a:=r_{i_{2}j_{2}}=1$ and $b:={\rm dim}(W^{\prime})$. Since $r=r_{11}+a+b\geq(a+1)+a+b$, we have $12a+b\leq r-$ 1. If the $S^{d}(W_{11}^{*})$-component $F_{11}$ of $F$ is zero, then we have
$F\in\sum_{1\leq d_{1}\leq d}S^{d_{1}}(W_{i_{2}j_{2}}^{*})\otimes S^{d-d_{1}}(W_{11}^{*}\oplus W^{\prime*})+\sum_{2\leq d_{2}\leq d}S^{d_{2}}(W^{\prime*})\otimes S^{d-d_{2}}(W_{11}^{*}\oplus W_{i_{2}j_{2}}^{*}),$
which contradicts the smoothness of $F$ by [OY19, Lemma 3.2]. Therefore, $F_{11}\neq 0$. Since the $S^{d-1}(W_{11}^{*})\otimes S^{1}(W_{i_{2}j_{2}}^{*})$-component $F_{11i_{2}j_{2}}$ of $F$ is not zero, we write $F_{11i_{2}j_{2}}=F_{11}^{\prime}\cdot L_{i_{2}j_{2}}^{\prime}$, where $F_{11}^{\prime}\in S^{d-1}(W_{11}^{*})$ and $L_{i_{2}j_{2}}^{\prime}\in S^{1}(W_{i_{2}j_{2}}^{*})$. By $\mathcal{A}=\{(i_{2},j_{2})\}$, we have $W_{11}\oplus W_{r_{r_{2}j_{2}}}$ is ${\rm Stab}_{G}(W_{11})$-stable. Let $M:=\{(g_{1},g_{2})|\;g_{1}=g|W_{11},g_{2}=g|W_{i_{2}j_{2}},g\in{\rm Stab}_{G}(W_{11})\}$. If $A=(\lambda_{1}I_{r_{11}},\lambda_{2})\in M$, then by $A(F)=F$, we have $\lambda_{1}^{d}=\lambda_{1}^{d-1}\lambda_{2}=1$ and $\lambda_{1}=\lambda_{2}$. Consider the surjective homomorphisms $\varphi_{1}:M\longrightarrow H_{11}$, $(g_{1},g_{2})\mapsto[g_{1}]$ and $\varphi_{2}:M\longrightarrow\widetilde{H}_{i_{2}j_{2}}\cong C_{d}$, $(g_{1},g_{2})\mapsto g_{2}$. For any $(g_{1},1)\in{\rm Ker}(\varphi_{2})$, we have $g_{1}$ preserves both $F_{11}$ and $F_{11i_{2}j_{2}}$, which implies that $\varphi_{1}|{\rm Ker}(\varphi_{2}):{\rm Ker}(\varphi_{2})\rightarrow H_{11}$ is injective. Clearly $|H_{11}|=|M|/d=|{\rm Ker}(\varphi_{2})|$. Thus $\{g_{1}|\;(g_{1},1)\in{\rm Ker}(\varphi_{2})\}\cong H_{11}$ is a desired $F_{11}$-lifting of $H_{11}$. This completes the proof of the lemma. ∎
6. Primitive groups of automorphisms
Finite linear groups in small dimensions have been completely classified (see e.g. [Fei70] for a survey). In particular, all finite primitive subgroups in ${\rm PGL}(r,{\mathbb{C}})$ for small $r$ are known. Based on this, we derive the following theorem which is the main result of this section.
Theorem 6.1.
Fix integers $n\geq 1$, $d\geq 3$. Let $F=F(x_{1},\ldots,x_{n+2})$ be a smooth form of degree $d\geq 3$ with primitive ${\rm Aut}(F)$ and $|{\rm Aut}(F)|\geq d^{n+2}\cdot(n+2)!$. Then up to isomorphism, $F$ is as follows:
$(n,d)$ ${\rm Lin}(X_{F})$ $|{\rm Lin}(X_{F})|$ $F$
$(1,4)$ ${\rm PSL}(2,7)$ $168$ $x_{1}^{3}x_{2}+x_{2}^{3}x_{3}+x_{3}^{3}x_{1}$
$(1,6)$ $C_{3}^{2}\rtimes{\rm SL}(2,3)$ $216$ $x_{1}^{6}+x_{2}^{6}+x_{3}^{6}-10(x_{1}^{3}x_{2}^{3}+x_{2}^{3}x_{3}^{3}+x_{3}^{3}x_{1}^{3})$
$(1,6)$ $A_{6}$ $360$ $10x_{1}^{3}x_{2}^{3}+9(x_{1}^{5}+x_{2}^{5})x_{3}-45x_{1}^{2}x_{2}^{2}x_{3}^{2}-135x_{1}x_{2}x_{3}^{4}+27x_{3}^{6}$
$(2,4)$ $C_{2}^{4}.S_{5}$ $1920$ $x_{1}^{4}+x_{2}^{4}+x_{3}^{4}+x_{4}^{4}+12x_{1}x_{2}x_{3}x_{4}$
$(4,6)$ ${\rm PSU}(4,3).C_{2}$ $6531840$ $\sum\limits_{1\leq i\leq 6}x_{i}^{6}+\sum\limits_{1\leq i\neq j\leq 6}15x_{i}^{4}x_{j}^{2}-\sum\limits_{1\leq i<j<k\leq 6}30x_{i}^{2}x_{j}^{2}x_{k}^{2}+240\sqrt{-3}x_{1}x_{2}x_{3}x_{4}x_{5}x_{6}$
Besides the values of ${\rm JC}(r)$, our proof of Theorem 6.1 relies on explicit list of the finite primitive subgroups in ${\rm PGL}(r,{\mathbb{C}})$ for $r\in\{2,3,4,5,6,8\}$, see Blichfeldt [Bli17] for $r\leq 4$ (see also [CS19, Appendix A] for diagrams indicating inclusions between groups in the case $r=4$); Brauer [Bra67] for $r=5$; Lindsey [Lin71] for $r=6$; Feit [Fei76] for $r=8$. For ease of reference, we present a list of finite primitive groups $H$ in ${\rm PGL}(r,{\mathbb{C}})$ with $|H|\geq 3^{r-1}r!$ and $r\in\{2,3,4,5,6\}$ (see Table 1).
$r$ $|H|$ $H$ $r$ $|H|$ $H$ $r$ $|H|$ $H$ $r$ $|H|$ $H$
2 12 $A_{4}$ 3 360 $A_{6}$ 4 1920 $C_{2}^{4}.S_{5}$ 4 25920 $\rm PSP(4,3)$
2 24 $S_{4}$ 4 720 $A_{4}\times A_{5}$ 4 1920 $C_{2}^{4}.S_{5}$ 5 25920 $\rm PSP(4,3)$
2 60 $A_{5}$ 4 720 $S_{6}$ 4 2520 $A_{7}$ 6 604800 $\rm HaJ$
3 60 $A_{5}$ 4 960 $C_{2}^{4}\rtimes A_{5}$ 4 3600 $A_{5}^{2}$ 6 3265920 $\rm PSU(4,3)$
3 72 $C_{3}^{2}\rtimes Q_{8}$ 4 960 $C_{2}^{4}\rtimes A_{5}$ 4 5760 $C_{2}^{4}\rtimes A_{6}$ 6 6531840 ${\rm PSU}(4,3).C_{2}$
3 168 ${\rm PSL}(2,7)$ 4 1152 $S_{4}^{2}\rtimes C_{2}$ 4 7200 $A_{5}^{2}\rtimes C_{2}$
3 216 $C_{3}^{2}\rtimes{\rm SL}(2,3)$ 4 1440 $S_{4}\times A_{5}$ 4 11520 $C_{2}^{4}.S_{6}$
Table 1. Large finite primitive groups $H\subset{\rm PGL}(r,{\mathbb{C}})$
Before presenting the proof of the theorem, we first give two lemmas.
Lemma 6.2.
Let $\widetilde{H}\subset{\rm GL}(r,{\mathbb{C}})$ be a finite primitive group. Let $H:=\pi(\widetilde{H})$. If $\widetilde{H}\cong H$, then the following statements hold:
  1. (1)
    $H$ admits a faithful linear representation of degree $r$;
  2. (2)
    $H$ has no non-trivial solvable normal subgroups;
  3. (3)
    For each prime $p$, $H$ has no subgroups isomorphic to $C_{p}^{r}$.
Proof.
(1) is clear. By $\widetilde{H}\cong\pi(\widetilde{H})$, we have $|Z(\widetilde{H})|=1$. For (2), we suppose $N$ is a non-trivial solvable normal subgroup of $\widetilde{H}$. Since $N$ is solvable, $N$ has a non-trivial characteristic abelian subgroup, say $N_{1}$. Then $N_{1}$ is a normal abelian subgroup of $\widetilde{H}$, which implies $N_{1}\subset Z(\widetilde{H})$ by primitivity of $\widetilde{H}$, a contradiction. Consequently, the statement (2) is true. For (3), we suppose that $\widetilde{H}$ has a subgroup, say $H_{1}$, isomorphic to $C_{p}^{r}$ for some prime $p$. Since $H_{1}\subset{\rm GL}(r,{\mathbb{C}})$, we infer that $A:={\rm diag}(\xi_{p},\ldots,\xi_{p})\in H_{1}$. Thus, $A\in Z(\widetilde{H})$, a contradiction. Hence (3) is true. ∎
By a theorem of Schur, for a finite group $G$, there is a finite group $SC(G)$ such that $SC(G)$ is a central extension of the Schur multiplier $H^{2}(G,{\mathbb{C}}^{\times})$ by $G$ and every projective representation of $G$ can be lifted to $SC(G)$ (cf. [Rot95, Chapter 7]). We call $SC(G)$ a Schur cover group of $G$. If $G$ is perfect, then $SC(G)$ is perfect and unique up to isomorphism (cf. [Rot95, Corollary 11.12]).
Lemma 6.3.
Let $H\subset{\rm PGL}(r,{\mathbb{C}})$ be a finite perfect subgroup preserving a form $F=F(x_{1},\ldots,x_{r})$ of degree $d\geq 3$. Then the Schur cover group $SC(H)$ of $H$ admits a linear representation $\rho:SC(H)\rightarrow{\rm GL}(r,{\mathbb{C}})$ such that $\rho(SC(H))\subset{\rm Aut}(F)$ and $\pi(\rho(SC(H)))=H$. Moreover, if $(r,d)=(4,6)$ and $F$ is smooth, then $\rho(SC(H))$ has no elements $A$ of order $5$ with ${\rm tr}(A)=-1$.
Proof.
The group $SC(H)$ has a representation $\rho:SC(H)\rightarrow{\rm GL}(r,{\mathbb{C}})$ such that $\pi(\rho(SC(H)))=H$. Since $H$ preserves $F$, we get a character $\chi:SC(H)\rightarrow{\mathbb{C}}^{\times}$ satisfying $\rho(g)(F)=\chi(g)F$ for all $g\in SC(H)$. On the other hand, $SC(H)$ is perfect since $H$ is perfect, which implies $\chi$ is trivial. Thus, $\rho(SC(H))$ preserves $F$. For the last statement in the lemma, we suppose $A={\rm diag}(\xi_{5},\xi_{5}^{2},\xi_{5}^{3},\xi_{5}^{4})\in{\rm Aut}(F)$. Then by $A(F)=F$, we have $x_{1}^{5}x_{i}\notin F$ for any $1\leq i\leq 4$, which contradicts smoothness of $F$. ∎
We are now ready to prove the main result of this section.
Proof of Theorem 6.1.
The primitivity of ${\rm Aut}(F)$ implies the only primitive constituent $H$ of ${\rm Aut}(F)$ is $\pi({\rm Aut}(F))\cong{\rm Aut}(F)/Z({\rm Aut}(F))$, where $Z({\rm Aut}(F))\cong C_{d}$. Hence, we have $|{\rm Aut}(F)|=B({\rm Aut}(F))=d\cdot|H|$ and $|H|\geq d^{n+1}(n+2)!$. Since $|H|\leq{\rm JC}(n+2)$, it suffices to consider the cases where $(n,d)$ is one of the following: $(1,3)$, $(1,4)$, $(1,5)$, $(1,6)$, $(1,7)$, $(2,3)$, $(2,4)$, $(2,5)$, $(2,6)$, $(2,7)$, $(2,8)$, $(2,9)$, $(2,10)$, $(3,3)$, $(4,3)$, $(4,4)$, $(4,5)$, $(4,6)$, $(6,3)$.
For $(n,d)$ equal to $(1,4)$, $(1,5)$, $(1,6)$, and $(1,7)$, the results are already known, see [Pam13] and [Har19]; for $(n,d)=(1,6)$ and $|{\rm Lin}(X_{F})|=216$, see also [BB22]; for $(n,d)=(1,3)$, see e.g. [YYZ23, Lemma 3.11]. For $2\leq n\leq 4$, Fermat cubic $n$-fold has the largest possible order for the automorphism group among all smooth cubic $n$-folds (see [Seg42], [Hos97], [Dol12] for $n=2$; [WY20] for $n=3$; [LZ22], [YYZ23] for $n=4$). For the case $(n,d)=(2,4)$, $F$ with primitive ${\rm Lin}(X_{F})$ of order $\geq 4^{3}\cdot 4!=1536$ is unique (up to isomorphism) and $|{\rm Lin}(X_{F})|=1920$ (see [Bur55, Exercise 6, Chap. XVII], [AOT24, Theorem A]).
Cases $(n,d)=(2,5),(2,7),(2,9),(4,5),(6,3)$. Since $n+2$ and $d$ are coprime, by [GLM23, Theorem 3.5], there is a (primitive) subgroup $\widetilde{H}\subset{\rm Aut}(F)$ with $\widetilde{H}\cong\pi(\widetilde{H})=H$. Thus, $H$ satisfies Lemma 6.2 (1)-(3). However, no such $H$ exists by Table 1 and [Fei76, Theorem A].
For $(n,d)=(2,6)$, we may assume $|H|\geq 6^{3}\cdot 4!=5184$. We then need to consider $H$ to be one of the following groups: $A_{5}^{2}\rtimes C_{2}$, $C_{2}^{4}\rtimes A_{6}$, $C_{2}^{4}.S_{6}$, ${\rm PSP}(4,3).$ Suppose $H\cong A_{5}^{2}\rtimes C_{2}$. Then $H$ has a perfect subgroup $H_{1}\cong A_{5}^{2}$. The Schur cover group of $A_{5}$ (resp. $A_{5}^{2}$) is ${\rm SL}(2,5)$ (resp. ${\rm SL}(2,5)^{2}$) of order $120$ (resp. $14400$). From the character table of ${\rm SL}(2,5)$ (see [CCNPW]), we infer that for each irreducible representation $\rho:{\rm SL}(2,5)^{2}\rightarrow{\rm GL}(4,{\mathbb{C}})$, there is $A\in\rho({\rm SL}(2,5)^{2})$ with $A$ of order $5$ and ${\rm tr}(A)=-1$, which contradicts Lemma 6.3. Thus, $H\ncong A_{5}^{2}\rtimes C_{2}$. Similarly, we have $H\ncong C_{2}^{4}\rtimes A_{6}$, $C_{2}^{4}.S_{6}$, ${\rm PSP}(4,3)$ by choosing $H_{1}$ as $A_{6}$, $A_{6}$, ${\rm PSP}(4,3)$ respectively.
For $(n,d)=(2,8),(2,10)$, we may assume $|H|\geq 8^{3}\cdot 4!=12288$. Thus we only need to consider $H\cong{\rm PSP}(4,3)$, which preserves no forms of degree $d=8,10$ by the Molien’s series of its Schur cover (cf. [ST54, Page 287]), a contradiction by Lemma 6.3.
For $(n,d)=(4,4)$, since $|H|\geq 4^{5}\cdot 6!=737280$, we have $H\cong{\rm PSU}(4,3)$ or $H\cong{\rm PSU}(4,3).C_{2}$. As in the previous cases, these two groups can be ruled out by the Molien’s series of the Schur cover of ${\rm PSU}(4,3)$ (cf. [Tod50, Page 86]) and the fact ${\rm PSU}(4,3)<{\rm PSU}(4,3).C_{2}$.
For $(n,d)=(4,6)$, there is a unique form $F=F(x_{1},\ldots,x_{6})$ of degree $6$ with primitive $\pi({\rm Aut}(F))\cong{\rm PSU}(4,3).C_{2}$ of order $6531840$ ([Tod50]). Smoothness of $F$ can be verified by Magma ([BCP97]). ∎
Forms preserved by finite primitive groups in ${\rm PGL}(2,{\mathbb{C}})$ are known (see [MBD16, §105]).
Lemma 6.4.
Let $H\subset{\rm PGL}(2,{\mathbb{C}})$ be a finite primitive group preserving a form $F=F(x_{1},x_{2})$ of degree $d\geq 1$. If $H=A_{4}$ (resp. $S_{4}$, resp. $A_{5}$), then $d\geq 4$ (resp. $d\geq 6$, resp. $d=12$ or $d\geq 20$). Moreover, if $(H,d)=(A_{4},4)$ (resp. $(S_{4},6)$, resp. $(A_{5},12)$), then $F$ is isomorphic to the smooth form $x_{1}^{4}+2\sqrt{-3}x_{1}^{2}x_{2}^{2}+x_{2}^{4}$ (resp. $x_{1}^{5}x_{2}+x_{2}^{5}x_{1}$, resp. $x_{1}^{11}x_{2}+11x_{1}^{6}x_{2}^{6}-x_{1}x_{2}^{11}$).
7. Proof of the main theorem
In this section, using Theorem 6.1 and the refined bounds in Section 5, we classify smooth forms with large imprimitive or reducible automorphism groups (Theorems 7.1 and 7.4). Combining these results, we prove a stronger form (Theorem 7.5) of Theorem 1.1.
For a smooth form $F$ of degree $d\geq 3$ in $r\geq 1$ variables, we define $\mathfrak{R}_{F}:=|{\rm Aut}(F)|/(d^{r}\cdot r!)$.
7.1. Imprimitive ${\rm Aut}(F)$
Recall that by an imprimitive finite linear group we mean that it is irreducible and not primitive. Now we classify smooth forms with large imprimitive automorphism groups.
Theorem 7.1.
Let $F=F(x_{1},\ldots,x_{n+2})$ be a smooth form of degree $d$, where $n\geq 1$, $d\geq 3$. Suppose ${\rm Aut}(F)$ is imprimitive and $F$ is not isomorphic to Fermat form $F_{d}^{n}$. If $|{\rm Aut}(F)|\geq d^{n+2}\cdot(n+2)!$, then up to isomorphism, $F$ is as follows:
$(n,d)$ ${\rm Aut}(X_{F})$ $|{\rm Aut}(X_{F})|$ $F$
$(2,6)$ $C_{6}.(S_{4}^{2}\rtimes C_{2})$ $6912$ $x_{1}^{5}x_{2}+x_{2}^{5}x_{1}+x_{3}^{5}x_{4}+x_{4}^{5}x_{3}$
$(2,12)$ $C_{12}.(A_{5}^{2}\rtimes C_{2})$ $86400$ $x_{1}^{11}x_{2}+11x_{1}^{6}x_{2}^{6}-x_{1}x_{2}^{11}+x_{3}^{11}x_{4}+11x_{3}^{6}x_{4}^{6}-x_{3}x_{4}^{11}$
$(4,12)$ $C_{12}^{2}.(A_{5}^{3}\rtimes S_{3})$ $186624000$ $x_{1}^{11}x_{2}+11x_{1}^{6}x_{2}^{6}-x_{1}x_{2}^{11}+x_{3}^{11}x_{4}+11x_{3}^{6}x_{4}^{6}-x_{3}x_{4}^{11}+x_{5}^{11}x_{6}+11x_{5}^{6}x_{6}^{6}-x_{5}x_{6}^{11}$
Proof.
By the assumption, $l({\rm Aut}(F))$ is of exponential type $r^{k}$ with $r\geq 2$ and $k\geq 2$, and ${\rm Aut}(F)$ has a primitive-constituent sequence $\mathcal{H}:=(H_{1},\ldots,H_{k})$ with $H_{i}\subset{\rm PGL}(r,{\mathbb{C}})$ and $H_{i}\cong H_{j}$ for $1\leq i,j\leq k$. By Lemma 4.6, we have $R(\mathcal{H},d)\geq 1$. By direct calculation (see (4.1)), we get $|H_{i}|>d^{r-1}\cdot r!$. It suffices to consider two cases: $r\geq 3$ and $r=2$.
Case $r\geq 3$. If $H_{1}$ preserves a smooth form $F_{1}(x_{1},\ldots,x_{r})$ of degree $d$, then by Theorem 6.1, we have $(r,d,|H_{1}|)=(3,4,168)$, $(3,6,360)$, $(4,4,1920)$, $(6,6,6531840)$. By direct calculation, we have $R(\mathcal{H},d)<1$, a contradiction. Thus, $H_{1}$ preserves no smooth forms of degree $d$. Since ${\rm JC}(r)>106$ for $r\geq 3$, by Lemmas 5.1, 5.5 and 4.14, we get $|{\rm Aut}(F)|<d^{n+2}\cdot(n+2)!$, a contradiction.
Case $r=2$. If $H_{1}$ preserves no smooth forms of degree $d$, then $|{\rm Aut}(F)|<d^{n+2}\cdot(n+2)!$ by similar arguments as in case $r\geq 3$. Thus, we may assume $H_{1}$ preserves a smooth form $F_{1}(x_{1},\ldots,x_{r})$ of degree $d$. By Lemma 6.4 , if $H_{1}\cong A_{4},S_{4},A_{5}$, then $H_{1}$ preserves no smooth forms in $2$ variables of degree less than $4,6,12$ respectively. If $H_{1}\cong A_{4}$, then $d\geq 4$ and $R(\mathcal{H},d)<1$, a contradiction. If $H_{1}\cong S_{4}$, then $d\geq 6$ and $R(\mathcal{H},d)\geq 1$ implies $(k,d)=(2,6)$. Then $R(\mathcal{H},d)=4/3$ and by Lemma 5.6, $F$ can be expressed as $F_{1}(x_{1},x_{2})+F_{2}(x_{3},x_{4})$, where $F_{i}$ are preserved by $H_{i}$. Therefore, $F$ is unique up to isomorphism by Lemma 6.4. If $H_{1}\cong A_{5}$, then $d=12$ or $d\geq 20$ by Lemma 6.4. So $R(\mathcal{H},d)\geq 1$ implies $(k,d)=(2,12),(3,12)$. Similar to the case $H_{1}=S_{4}$, we have the uniqueness of $F$. ∎
As a consequence of Theorems 6.1, 7.1 and Lemma 6.4, we have the following result.
Proposition 7.2.
Let $F=F(x_{1},\ldots,x_{r})$ be a smooth form of degree $d\geq 3$, where $r\geq 2$. Suppose ${\rm Aut}(F)$ is irreducible and $\mathfrak{R}_{F}=\frac{|{\rm Aut}(F)|}{d^{r}\cdot r!}>1$. Then either $\mathfrak{R}_{F}=5/2$ or $\mathfrak{R}_{F}\leq 25/12$. Moreover, $\mathfrak{R}_{F}=5/2$ if and only if $(r,d)=(2,12)$ and ${\rm Aut}(F)/Z({\rm Aut}(F))\cong A_{5}$.
7.2. Reducible ${\rm Aut}(F)$
The following lemma will be used to prove Theorem 7.4.
Lemma 7.3.
Let $n_{1},\ldots,n_{m}$ be positive integers, where $m\geq 2$. Suppose $q_{i}$ $(1\leq i\leq m)$ are rational numbers satisfying $q_{i}=5/2$ (resp. $q_{i}=1$) if $n_{i}\geq 2$ (resp. $n_{i}=1$). If $\frac{\prod_{i=1}^{m}q_{i}\cdot n_{i}!}{(n_{1}+\cdots+n_{m})!}\geq 1$, then $m=2$ and $n_{1}=n_{2}=2$.
Proof.
We may assume $n_{1}\geq n_{2}\geq\cdots\geq n_{m}$. We write
$q:=\frac{\prod_{i=1}^{m}q_{i}\cdot n_{i}!}{(n_{1}+\cdots+n_{m})!}=\frac{q_{1}\cdot n_{1}!n_{2}!}{(n_{1}+n_{2})!}\cdot\frac{q_{2}\cdot(n_{1}+n_{2})!n_{3}!}{(n_{1}+n_{2}+n_{3})!}\cdots\frac{q_{m-1}q_{m}(n_{1}+\cdots+n_{m-1})!n_{m}!}{(n_{1}+\cdots+n_{m})!}.$
By direct calculations, if $n_{m}=1$ or $m\geq 3$ or $n_{1}>2$, then all fractions on the right side are less than $1$, a contradiction to $q\geq 1$. Thus, we have $m=2$, $n_{1}=n_{2}=2$ and $q=\frac{25}{24}$. ∎
Next we show that smooth homogeneous polynomials of reducible automorphism groups have fewer automorphisms than Fermat polynomials.
Theorem 7.4.
Let $F=F(x_{1},\ldots,x_{r})$ be a smooth form of degree $d\geq 3$, where $r\geq 3$. If ${\rm Aut}(F)$ is reducible, then $|{\rm Aut}(F)|<d^{r}\cdot r!$
Proof.
To prove the theorem by contradiction, we suppose $\mathfrak{R}_{F}=\frac{|{\rm Aut}(F)|}{d^{r}\cdot r!}\geq 1$. Let $G:={\rm Aut}(F)$. By Theorem 3.9, we may assume at least one subdegree of $G$ is larger than $1$. Let $V_{i},W_{ij},k_{i},r_{ij},H_{ij},m,F_{ij}$ be as in $\bf{Set\textrm{-}up}$ 5.3. We define $\beta$ be to $1$ (resp. $2$) if $l(G)$ has at least two different subdegrees (resp. all subdegrees of $G$ are the same). From the definitions, we have
$\frac{B(G,F)}{|H_{ij}|^{k}}\leq\frac{R(l(G),d)\cdot d^{r}\cdot r!}{\beta\cdot{\rm JC}(r_{ij})^{k}}$ (7.1)
for any pair $(i,j)$ and any $1\leq k\leq k_{i}$. By Lemmas 4.6 and 4.14, we obtain $R(l(G),d)<60\beta$. If there is a monomial
$\mathfrak{m}\in S^{1}(W_{i_{1}j_{1}}^{*})\otimes S^{d-1}(\oplus_{r_{i_{2}j_{2}}=1}W_{i_{2}j_{2}}^{*})$
in $F$ with $r_{i_{1}j_{1}}\geq 2$, then by Lemma 5.4, we have
$|{\rm Aut}(F)|\leq\frac{B(G,F)}{|H_{i_{1}j_{1}}|^{k_{i_{1}}}}\leq\frac{R(l(G),d)\cdot d^{r}\cdot r!}{\beta\cdot{\rm JC}(r_{i_{1}j_{1}})^{k_{i_{1}}}}<d^{r}\cdot r!,$
a contradiction. From now on, we may assume that such monomial $\mathfrak{m}$ does not exist. Without loss of generality, we assume $r_{ij}\geq r_{i^{\prime}j^{\prime}}$ whenever $i<i^{\prime}$. Then there exists a positive integer $m_{1}\leq m$ such that $r_{ij}\geq 2$ for $i\leq m_{1}$ and $r_{ij}=1$ for $i>m_{1}$. We set
$F_{i}:=\sum_{1\leq j\leq k_{i}}F_{ij},\;(i=1,2,\ldots,m_{1})~{}~{}~{}{\rm and}~{}~{}~{}F_{m_{1}+1}:=\sum_{\sum_{i>m_{1},1\leq j\leq k_{i}}d_{ij}=d}F^{(d_{11},d_{12},\ldots,d_{mk_{m}})}.$
If $m_{1}=m$, by definition $F_{m_{1}+1}=0$. Then Lemma 5.1 implies that $F_{m_{1}+1}$ is smooth if $m_{1}<m$. It suffices to consider two cases: (1) $F_{ij}$ are smooth for all pairs $(i,j)$ with $i\leq m_{1}$; (2) $F_{i_{1}j_{1}}$ is not smooth for some pair $(i_{1},j_{1})$ with $i_{1}\leq m_{1}$.
Case (1). Note that $V_{i}$ $(i=1,\ldots,m_{1})$ are irreducible $G$-stable subspaces of dimension $r_{i}:={\rm dim}(V_{i})=r_{i1}k_{i}$ and the irreducible linear groups $G_{i}:=p_{i}(G)$ preserve the smooth forms $F_{i}$ of degree $d$. Here $p_{i}:G\rightarrow{\rm GL}(V_{i})$ is given by $g\mapsto g|V_{i}$. Let $V^{\prime}:=\oplus_{i>m_{1}}V_{i}$ and $p^{\prime}:G\rightarrow{\rm GL}(V^{\prime})$ be given by $g\mapsto g|V^{\prime}$. Let $G_{m_{1}+1}:=p^{\prime}(G)$. Then $r=r_{1}+\cdots r_{m_{1}}+r_{m_{1}+1}$, where $r_{m_{1}+1}:={\rm dim}(V^{\prime})$. By Proposition 7.2, we have $\mathfrak{R}_{F_{i}}=\frac{|G_{F_{i}}|}{d^{r_{i}}\cdot r_{i}!}\leq 5/2$ $(i=1,2,\ldots,m_{1})$. By Theorem 3.9, we have $\mathfrak{R}_{F_{m_{1}+1}}=\frac{|G_{F_{m_{1}+1}}|}{d^{r_{m_{1}+1}}\cdot r_{m_{1}+1}!}\leq 1$. Clearly $G_{i}\subset G_{F_{i}}$ $(i=1,2,\ldots,m_{1}+1)$. Then by Lemma 7.3 and
$1\leq\mathfrak{R}_{F}=\frac{|{\rm Aut}(F)|}{d^{r}\cdot r!}\leq\frac{|G_{1}|\cdot|G_{2}|\cdots|G_{m_{1}}|\cdot|G_{m_{1}+1}|}{d^{r}\cdot r!}\leq\frac{\prod_{i=1}^{m_{1}+1}\mathfrak{R}_{F_{i}}\cdot r_{i}!}{(r_{1}+\cdots+r_{m_{1}+1})!},$
we have $m_{1}=m=2$, $r_{1}=r_{2}=2$ and $r=4$. Thus by Proposition 7.2 again, we infer that $H_{11}\cong H_{21}\cong A_{5}$ and $d=12$. Then we may write $F=F_{1}(x_{1},x_{2})+F_{2}(x_{3},x_{4})+F^{\prime}(x_{1},x_{2},x_{3},x_{4})$, where
$F^{\prime}\in\mathop{\vbox{\hbox{{$\bigoplus$}}}}_{1\leq d_{1}\leq 11}S^{d_{1}}(V_{1}^{*})\otimes S^{12-d_{1}}(V_{2}^{*}).$
If $F^{\prime}$ is zero, then ${\rm Aut}(F)$ is irreducible (see Theorem 7.1), a contradiction. If $F^{\prime}$ is not zero, then by Lemmas 5.6 and 6.4, we get $\mathfrak{R}_{F}<1$, a contradiction. This proves case (1).
Case (2). Let $\mathcal{A}$ and $c$ be as in Lemma 5.7. Then $c>0$. If there is a pair $(i_{2},j_{2})\in\mathcal{A}$ with $r_{i_{2}j_{2}}\geq 2$, then $\mathfrak{R}_{F}<1$ by Lemma 5.7 (1) and the inequality (7.1). So we may assume $r_{i_{2},j_{2}}=1$ for all $(i_{2},j_{2})\in\mathcal{A}$. If $c\geq 2$, then by Lemmas 4.14 and 5.7, we have $\mathfrak{R}_{F}<1$, which is impossible. Thus we have $c=1$ and $H_{i_{1}j_{1}}$ admits an $F_{i_{1}j_{1}}$-lifting preserving some non-zero form $F_{1}^{\prime}$ of degree $d-1$. Since $l({\rm Aut}(F))$ contains $r_{i_{1}j_{1}}\geq 2$ and $1$, by Lemma 4.14 and Remark 4.15, we have $r_{i_{1}j_{1}}=2$, $3$, $4$, $6$ and $R(l(G),d)<11$, $2$, $11$, $6$ respectively. By Lemma 4.6, we have
$1\leq\mathfrak{R}_{F}\leq\frac{R(l(G),d)\cdot|H_{r_{i_{1}j_{1}}}|}{{\rm JC}(r_{i_{1}j_{1}})}\;\text{ and }\;|H_{r_{i_{1}j_{1}}}|\geq\frac{{\rm JC}(r_{i_{1}j_{1}})}{R(l(G),d)}.$
From this, we have $|H_{i_{1}j_{1}}|>60/11,360/2,25920/11,6531840/6$ if $r_{i_{1}j_{1}}=2$, $3$, $4$, $6$ respectively. On the other hand, $H_{i_{1}j_{1}}$ admitting an $F_{i_{1}j_{1}}$-lifting implies that $H_{i_{1}j_{1}}$ satisfies Lemma 6.2 (1)-(3), which is impossible by Table 1 (see e.g. [CCNPW] for character tables of simple groups). ∎
7.3. A stronger form of the main theorem
Now we prove the following theorem, which in particular completes the proof of Theorem 1.1.
Theorem 7.5.
Fix integers $n\geq 1$, $d\geq 3$ with $(n,d)\neq(1,3),(2,4)$. Let $X\subset{\mathbb{P}}^{n+1}$ be a smooth hypersurface of degree $d$ with $|{\rm Aut}(X)|\geq d^{n+1}\cdot(n+2)!$. If $X$ is not isomorphic to Fermat hypersurface $X_{d}^{n}$, then $X$ is isomorphic to one of the following smooth hypersurfaces $X_{F}$:
$(n,d)$ ${\rm Aut}(X_{F})$ $|{\rm Aut}(X_{F})|$ $F$
$(1,4)$ ${\rm PSL}(2,7)$ $168$ $x_{1}^{3}x_{2}+x_{2}^{3}x_{3}+x_{3}^{3}x_{1}$
$(1,6)$ $C_{3}^{2}\rtimes{\rm SL}(2,3)$ $216$ $x_{1}^{6}+x_{2}^{6}+x_{3}^{6}-10(x_{1}^{3}x_{2}^{3}+x_{2}^{3}x_{3}^{3}+x_{3}^{3}x_{1}^{3})$
$(1,6)$ $A_{6}$ $360$ $10x_{1}^{3}x_{2}^{3}+9(x_{1}^{5}+x_{2}^{5})x_{3}-45x_{1}^{2}x_{2}^{2}x_{3}^{2}-135x_{1}x_{2}x_{3}^{4}+27x_{3}^{6}$
$(2,6)$ $C_{6}.(S_{4}^{2}\rtimes C_{2})$ $6912$ $x_{1}^{5}x_{2}+x_{2}^{5}x_{1}+x_{3}^{5}x_{4}+x_{4}^{5}x_{3}$
$(2,12)$ $C_{12}.(A_{5}^{2}\rtimes C_{2})$ $86400$ $x_{1}^{11}x_{2}+11x_{1}^{6}x_{2}^{6}-x_{1}x_{2}^{11}+x_{3}^{11}x_{4}+11x_{3}^{6}x_{4}^{6}-x_{3}x_{4}^{11}$
$(4,6)$ ${\rm PSU}(4,3).C_{2}$ $6531840$ $\sum\limits_{1\leq i\leq 6}x_{i}^{6}+\sum\limits_{1\leq i\neq j\leq 6}15x_{i}^{4}x_{j}^{2}-\sum\limits_{1\leq i<j<k\leq 6}30x_{i}^{2}x_{j}^{2}x_{k}^{2}+240\sqrt{-3}x_{1}x_{2}x_{3}x_{4}x_{5}x_{6}$
$(4,12)$ $C_{12}^{2}.(A_{5}^{3}\rtimes S_{3})$ $186624000$ $x_{1}^{11}x_{2}+11x_{1}^{6}x_{2}^{6}-x_{1}x_{2}^{11}+x_{3}^{11}x_{4}+11x_{3}^{6}x_{4}^{6}-x_{3}x_{4}^{11}+x_{5}^{11}x_{6}+11x_{5}^{6}x_{6}^{6}-x_{5}x_{6}^{11}$
Proof.
Let $F_{0}$ be a defining polynomial of $X$. Based on the exact sequence (1.1), by [MM63] and [Cha78], we have $|{\rm Aut}(X)|=|{\rm Lin}(X)|=|{\rm Aut}(F_{0})|/d$. Thus, $|{\rm Aut}(F_{0})|\geq d^{n+2}\cdot(n+2)!$, which implies ${\rm Aut}(F_{0})$ is irreducible by Theorem 7.4. Then the conclusion of the theorem follows from Theorems 3.9, 6.1, 7.1. ∎
Remark 7.6.
Let $X_{F}\subset{\mathbb{P}}^{n+1}$ be a smooth hypersurface of degree $d\geq 3$ defined by $F$. If $n+2\geq 71$, then ${\rm J}(n+2)=(n+3)!$ by [Col07, Theorem A]. On the other hand, by Theorem 3.4, the optimal upper bound for the size of an abelian subgroup in ${\rm Aut}(F)$ is $d^{n+2}$. Thus the product ${\rm J}(n+2)\cdot d^{n+2}=d^{n+2}\cdot(n+3)!$ gives an upper bound for $|{\rm Aut}(F)|$, which gives an upper bound $d^{n+1}\cdot(n+3)!$ for $|{\rm Aut}(X_{F})|$. However, this bound for $|{\rm Aut}(X_{F})|$ is larger than the optimal bound $d^{n+1}\cdot(n+2)!=|{\rm Aut}(X_{d}^{n})|$.
Remark 7.7.
A smooth hypersurface $X$ in ${\mathbb{P}}^{n+1}$ of degree $d=1$ (resp. $2$) is isomorphic to ${\mathbb{P}}^{n}$ (resp. Fermat quadric $X_{2}^{n}$) with infinite automorphism group. Smooth cubic curves are elliptic curves and their automorphism groups are infinite. The automorphism group of a generic quartic surface is finite, but there are many smooth quartic surfaces with infinite automorphism groups. The automorphism group of Fermat quartic $X_{4}^{2}$ is infinite, and its finite subgroup ${\rm Lin}(X_{4}^{2})\cong C_{4}^{3}\rtimes S_{4}$ determines $X_{4}^{2}$ uniquely among all K3 surfaces ([Ogu05]).